Jefferson Lab HALL A Logbook for October 10 - All EXCEPT Automatic Entries

  • Main INDEX

    Sunday

  • 337892 10/31/10 23:19 A. Tobias Other ROCs for run 8187
  • 337891 10/31/10 23:18 A. Tobias RunControl and Scalars for run 8187
  • 337890 10/31/10 23:17 A. Camsonne Moved clear of VME ADCs in DVCS crates to prestart
  • 337889 10/31/10 23:17 A. Tobias ROCs for run 8187
  • 337886 10/31/10 22:34 A. Tobias Other ROCs for run 8186
  • 337885 10/31/10 22:33 A. Tobias RunControl and Scalars for run 8186
  • 337884 10/31/10 22:32 A. Tobias ROCs for run 8186
  • 337883 10/31/10 22:02 Abdurahim Target screen & Strip charts
  • 337881 10/31/10 21:53 Abdurahim Compton Vertical Scan
  • 337880 10/31/10 21:53 A. Tobias RuncControl for Run 8185
  • 337878 10/31/10 21:53 A. Camsonne Clock with a large prescale for data taking
  • 337877 10/31/10 21:53 A. Tobias ROCs and Scalars for run 8185
  • 337875 10/31/10 21:35 A. Camsonne Random trigger on for DVCS at about 11 Hz
  • 337874 10/31/10 21:29 Ramesh DVCS calo anode current
  • 337872 10/31/10 21:25 A. Tobias Other ROCs for run 8184
  • 337871 10/31/10 21:25 A. Tobias RunControl and Scalars for run 8184
  • 337870 10/31/10 21:23 A. Tobias RunControl and Scalars for run 8184
  • 337869 10/31/10 21:22 A. Tobias ROCs for run 8184
  • 337866 10/31/10 20:38 A. Tobias Other ROCs for run 8183
  • 337865 10/31/10 20:36 A. Tobias RunControl and Scalars for run 8183
  • 337864 10/31/10 20:35 A. Tobias ROCs for run 8183
  • 337856 10/31/10 19:46 A. Tobias Other ROCs for 8182
  • 337855 10/31/10 19:45 A. Tobias RunControl and Scalars for run 8182
  • 337854 10/31/10 19:44 A. Tobias ROCs for run 8182
  • 337853 10/31/10 19:12 Abdurahim Target screen & Strip charts
  • 337850 10/31/10 19:01 A. Tobias Other ROCs for run 8181
  • 337849 10/31/10 19:00 A. Tobias RunControl & Scalars for run 8181
  • 337847 10/31/10 18:59 A. Tobias ROCs for run 8181
  • 337841 10/31/10 18:07 A. Tobias Other ROCs for run 8180
  • 337840 10/31/10 18:06 A. Tobias RunControl & Scalars for run 8180
  • 337839 10/31/10 18:05 A. Tobias ROCs for run 8180
  • 337837 10/31/10 17:23 Abdurahim Compton cavity lock script has been changed
  • 337836 10/31/10 17:10 Abdurahim Target -> LH2, Loop2, 15 cm
  • 337835 10/31/10 16:56 A. Tobias RunControl for run 8179
  • 337833 10/31/10 16:53 A. Tobias ROCs for run 8179
  • 337832 10/31/10 16:50 Ramesh Target moved to empty one
  • 337831 10/31/10 16:46 A. Tobias TS11 still has ERROR in predicted helicity. You are losing data !!
  • 337827 10/31/10 16:21 A. Tobias Other ROCs for run 8176
  • 337826 10/31/10 16:19 A. Tobias RunControl & Scalars for Run 8176
  • 337825 10/31/10 16:18 A. Tobias ROCs for run 8176
  • 337824 10/31/10 16:17 A. Tobias TS11 Errors run 8176
  • 337821 10/31/10 16:08 A. Tobias Other ROCs for run 8175
  • 337820 10/31/10 16:07 A. Tobias RunControl & Scalars for Run 8175
  • 337819 10/31/10 16:06 A. Tobias ROCs for run 8175
  • 337817 10/31/10 15:59 bogdanw shift summary
  • 337812 10/31/10 14:19 chench ROCs for run8173
  • 337811 10/31/10 14:18 chench DAQ and scalers for run8173
  • 337808 10/31/10 13:35 chench ROCs for run 8172
  • 337807 10/31/10 13:34 chench DAQ and scalers for run8172
  • 337804 10/31/10 12:45 chench DAQ and scalers for run8171
  • 337803 10/31/10 12:44 chench ROCs for run8171
  • 337800 10/31/10 11:59 chench DAQ and scalers for run8170
  • 337799 10/31/10 11:57 A. Camsonne Compton rates flatlined at 0 - rebooting the Compton DAQ
  • 337798 10/31/10 11:56 chench Rocs for run8170
  • 337785 10/31/10 10:53 pzhu PbF2 Calorimeter ADC Readback
  • 337784 10/31/10 10:50 chench DAQ and scalers for run8169
  • 337783 10/31/10 10:47 chench ROCs for run(8169)
  • 337780 10/31/10 09:50 chench DAQ and scalers (run8618)
  • 337779 10/31/10 09:49 chench Rocs for run 8168
  • 337773 10/31/10 08:09 hyde Cerenkov efficiency
  • 337772 10/31/10 07:48 hyde Owl Shift Summary
  • 337769 10/31/10 07:40 D.Wang rocs (8165)
  • 337768 10/31/10 07:39 D.Wang coda and scalers (8165)
  • 337765 10/31/10 06:52 D.Wang rocs (8164)
  • 337764 10/31/10 06:52 hyde Normalization check
  • 337763 10/31/10 06:51 D.Wang coda and scalers(8164)
  • 337761 10/31/10 05:59 lselvy Target Update
  • 337758 10/31/10 05:19 D.Wang rocs (8162)
  • 337757 10/31/10 05:18 D.Wang coda and scalers (8162)
  • 337756 10/31/10 04:48 hyde coda crash
  • 337755 10/31/10 04:45 hyde Normalization Runs
  • 337751 10/31/10 04:02 lselvy Helicity Error
  • 337743 10/31/10 03:51 D.Wang rocs (8159)
  • 337742 10/31/10 03:51 D.Wang coda and scalers (8159)
  • 337741 10/31/10 03:47 hyde calo dc anode current
  • 337740 10/31/10 03:45 hyde beam strip chart
  • 337739 10/31/10 03:41 hyde Cerenkov timing in coincidence
  • 337738 10/31/10 03:34 hyde Hall A Tools Snapshot
  • 337735 10/31/10 03:05 D.Wang rocs (8158)
  • 337734 10/31/10 03:04 D.Wang coda and scalers(8158)
  • 337726 10/31/10 02:16 Carlos Production in buffering mode
  • 337725 10/31/10 02:12 D.Wang rocs (8157)
  • 337724 10/31/10 02:12 D.Wang coda and scalers (8157)
  • 337723 10/31/10 02:07 HYDE DVCS CALO Anode current
  • 337711 10/31/10 00:58 hyde raster
  • 337709 10/31/10 00:42 hyde calo HV on
  • 337706 10/31/10 00:28 hyde calo HV off
  • 337700 10/31/10 00:01 A. Camsonne End of run 8144
  • Saturday

  • 337698 10/30/10 23:58 A. Tobias Target Status: Loop 2 LH2
  • 337697 10/30/10 23:54 A. Tobias Target Stripcharts
  • 337696 10/30/10 23:53 Ramesh Shift Summary
  • 337694 10/30/10 23:51 Abdurahim ROCs and Scaler screen snapshot
  • 337692 10/30/10 23:43 A. Tobias Loop 2 (LH2) Fan set to 19% = 46Hz
  • 337690 10/30/10 23:30 A. Camsonne Run with buffering 8143
  • 337687 10/30/10 23:17 Ramesh ps9=1 run for 10 min
  • 337684 10/30/10 23:14 Abdurahim ROCs and Scaler run8140
  • 337683 10/30/10 23:13 A. Camsonne Compton HV on - started run in backroom -
  • 337681 10/30/10 22:18 Abdurahim TS11 is fine now
  • 337679 10/30/10 22:15 Abdurahim Error massage in TS11
  • 337675 10/30/10 21:38 R. Michaels re: number of events in halog end-of-run
  • 337674 10/30/10 21:25 Abdurahim ROCs and Scalers snapshot run8137
  • 337671 10/30/10 20:36 Abdurahim ROC s and Scalers before the end of run8136
  • 337668 10/30/10 20:19 A. Tobias Target Screenshots
  • 337661 10/30/10 19:27 Abdurahim ROCs, Scalers and CODA screen snapshots before the end of 8135
  • 337659 10/30/10 19:21 Ramesh Discrepancy between Coda_event_number and end_of_run_event_number
  • 337658 10/30/10 19:11 Abdurahim CALO TDC histo mismacth run8133
  • 337657 10/30/10 19:09 Abdurahim BPMAx histo mismatch run8133
  • 337656 10/30/10 19:05 A. Tobias Target: Loop 2 Fan settings
  • 337655 10/30/10 19:00 A. Tobias Target: Loop 2 Fan Settings
  • 337654 10/30/10 19:00 Ramesh Calorimeter current
  • 337653 10/30/10 18:59 A. Tobias Loop 1 Fan speed recovers to 24Hz
  • 337647 10/30/10 18:55 Abdurahim mismatched histos in CALO TDC run8134
  • 337646 10/30/10 18:53 Abdurahim mismatched histo in BPM run8134
  • 337645 10/30/10 18:48 Abdurahim Strange Histo in Left arm S2 ADC
  • 337644 10/30/10 18:37 Ramesh Halog down intermittently
  • 337643 10/30/10 18:37 A. Tobias CryoTarget Fans
  • 337642 10/30/10 18:23 Ramesh Flaky Q3 liquid level!
  • 337641 10/30/10 18:23 Abdurahim ROCs, Scalers and CODA screens before the end of run8134
  • 337638 10/30/10 17:28 Abdurahim ROCS and CODA screen snapshot before the end of 8133
  • 337635 10/30/10 16:30 A. Tobias Target moved to Loop2 H2 15cm
  • 337633 10/30/10 16:14 A. Camsonne Added tsUtil.o to trigger_left.boot
  • 337631 10/30/10 16:03 bogdanw shift summary
  • 337629 10/30/10 15:19 A. Camsonne Sync ADC work
  • 337628 10/30/10 15:19 R. Michaels neg. deadtimes fixed
  • 337557 10/30/10 12:01 adaq hall status
  • 337547 10/30/10 11:29 A. Camsonne DVCS Trigger firmware loaded Run 8091
  • 337546 10/30/10 11:24 A. Camsonne Empty FIFO messages in ROC14 and ROC18
  • 337545 10/30/10 11:20 A. Camsonne DVCS Trigger module pedestals
  • 337535 10/30/10 10:19 flay Run 8089: S2m ADCs
  • 337533 10/30/10 10:15 flay Run 8090 Stopped Early
  • 337530 10/30/10 10:03 flay Run 8089 Trigger/ROC Snapshots
  • 337529 10/30/10 09:15 flay Run 8088: Lumis (BCM Normalized) Rates vs. Event # Has Odd Structure
  • 337526 10/30/10 09:01 flay Run 8088 Trigger/ROC Snapshots
  • 337525 10/30/10 08:26 Carlos Re: Compton history plot
  • 337524 10/30/10 08:23 adaq Compton CODA
  • 337522 10/30/10 08:19 efuchey 10-30-10 Owl Shift summary
  • 337520 10/30/10 08:01 efuchey Compton data taking restarted
  • 337518 10/30/10 07:51 D.Wang rocs (8087)
  • 337517 10/30/10 07:50 D.Wang coda and scalers (8087)
  • 337514 10/30/10 07:07 D.Wang rocs (8086)
  • 337513 10/30/10 07:07 D.Wang coda and scalers (8086)
  • 337512 10/30/10 06:46 efuchey Compton Rates 10-31
  • 337511 10/30/10 06:27 efuchey Rates going wild
  • 337508 10/30/10 06:08 lselvy Target Update
  • 337507 10/30/10 06:05 D.Wang rocs (8085)
  • 337506 10/30/10 06:04 D.Wang coda and scalers (8085)
  • 337504 10/30/10 05:13 D.Wang rocs (8084)
  • 337503 10/30/10 05:13 D.Wang coda and scalers (8084)
  • 337498 10/30/10 04:19 D.Wang rocs (8082)
  • 337497 10/30/10 04:18 D.Wang coda and scalers (8082)
  • 337489 10/30/10 03:29 D.Wang rocs (8081)
  • 337488 10/30/10 03:28 D.Wang coda and scalers (8081)
  • 337487 10/30/10 03:01 efuchey Update on the asymmetry history
  • 337481 10/30/10 02:42 efuchey First asymmetry history
  • 337478 10/30/10 02:34 D.Wang rocs (8080)
  • 337477 10/30/10 02:33 D.Wang coda and scalers (8080)
  • 337476 10/30/10 01:56 efuchey Junky Compton "production"
  • 337473 10/30/10 01:50 D.Wang rocs (8079)
  • 337472 10/30/10 01:49 D.Wang coda and scalers (8079)
  • 337469 10/30/10 00:59 D.Wang rocs (8078)
  • 337468 10/30/10 00:59 D.Wang coda and scalers (8078)
  • 337467 10/30/10 00:21 efuchey Hall A tools screenshot.
  • 337466 10/30/10 00:19 efuchey Calo base currents
  • 337463 10/30/10 00:10 D.Wang rocs (run 8077)
  • 337462 10/30/10 00:09 D.Wang coda and scalers
  • 337460 10/30/10 00:00 GOLGE swing shift summary
  • Friday

  • 337459 10/29/10 23:48 efuchey Compton rates after beam trip
  • 337458 10/29/10 23:24 Golge ARC magnet trip, beam off
  • 337453 10/29/10 22:11 chench screenshot of roc18 for run8075
  • 337452 10/29/10 22:10 chench screenshot for ROC14 and 17 for run8075
  • 337451 10/29/10 22:10 chench screenshot of DAQ window and scalers for run8075
  • 337448 10/29/10 21:17 chench correction: the run muber is 8074
  • 337447 10/29/10 21:16 chench corrections: the run number is 8074
  • 337446 10/29/10 21:14 chench screenshot of roc18 for run8074
  • 337445 10/29/10 21:13 chench screenshot of ROC14 and 17 for run8075
  • 337444 10/29/10 21:13 chench screenshot of DAQ window and scalers for run8075
  • 337441 10/29/10 20:22 chench screenshot of ROC18 for run8073
  • 337440 10/29/10 20:21 chench screenshot of ROC14 and ROC17
  • 337439 10/29/10 20:21 chench Screenshot of DAQ window and scalers for run8073
  • 337431 10/29/10 19:30 chench screenshot of roc18 for run8072
  • 337430 10/29/10 19:30 chench screenshot of ROC14 and 17 for run8072
  • 337429 10/29/10 19:29 chench screenshot of DAQ window and scalers for run8072
  • 337421 10/29/10 18:41 A. Camsonne Added clear of SIS scalers in DVCS crates
  • 337420 10/29/10 18:32 chench screenshot for roc18 for run8071
  • 337419 10/29/10 18:32 chench screenshot of ROC14 and ROC17 for run8071
  • 337418 10/29/10 18:31 chench Screenshot of DAQ window and scalers for run8071
  • 337416 10/29/10 17:41 chench Screenshot of roc18 for run8070
  • 337415 10/29/10 17:41 chench Sreenshot of ROC14 and ROC17 for run8070
  • 337414 10/29/10 17:40 chench Screenshot of DAQ window and scalers for run8070
  • 337413 10/29/10 17:32 hyde clock consistency check
  • 337410 10/29/10 16:54 chench screenshot of ROC18 for run8069
  • 337409 10/29/10 16:53 chench screenshot for ROC14 and ROC17 for run8069
  • 337408 10/29/10 16:52 chench Screenshot of DAQ and scalers for run 8069
  • 337407 10/29/10 16:22 efuchey Compton rates for 10-29-10
  • 337404 10/29/10 16:03 A. Camsonne DAQ Sync work
  • 337403 10/29/10 16:02 Saha Shift Summary
  • 337399 10/29/10 15:52 R. Michaels target to LH2, 15 cm
  • 337398 10/29/10 15:45 Ole Hansen Analyzer updated
  • 337397 10/29/10 15:38 R. Michaels moving to empty target
  • 337372 10/29/10 14:27 efuchey QWP scan for 10/29/10
  • 337360 10/29/10 13:55 hyde DVCS trig scaler
  • 337341 10/29/10 12:51 R. Michaels helicity stable for 15 min
  • 337340 10/29/10 12:39 R. Michaels occassional "bad" helicity
  • 337338 10/29/10 12:28 R. Michaels re: last scaler in datastream
  • 337337 10/29/10 12:28 A. Camsonne Sync ADC ROC3 and ROC14
  • 337336 10/29/10 12:25 A. Camsonne Sync ADC for ROC17 and ROC4 look ok - need to check ROC3 and ROC14
  • 337328 10/29/10 11:51 kelleher warming up Moller
  • 337327 10/29/10 11:35 Carlos Re: evtyebits problems
  • 337326 10/29/10 11:13 R. Michaels evtyebits problems.
  • 337323 10/29/10 11:00 Punjabi run 8035 ROC screens
  • 337322 10/29/10 10:59 Punjabi RC and scaler screens
  • 337316 10/29/10 10:55 jpchen target is fine
  • 337315 10/29/10 10:26 R. Michaels updated halla_scaler_process (halog scalers)
  • 337314 10/29/10 10:17 Carlos Deadtime study special runs
  • 337300 10/29/10 09:05 Punjabi End of run ROC screens
  • 337298 10/29/10 08:22 R. Michaels target instabilities
  • 337295 10/29/10 08:05 jroche owl shift summary
  • 337293 10/29/10 08:00 flay Run 8026 Trigger/ROC Snapshots
  • 337292 10/29/10 07:10 jroche "new" results for the BCM calibration
  • 337289 10/29/10 07:06 flay Run 8026 Trigger/ROC Snapshots
  • 337288 10/29/10 06:24 jroche Counting room somewhat cold since an hour or so
  • 337287 10/29/10 06:18 jroche End of run halog not showing up or very slow to show up
  • 337283 10/29/10 06:09 flay Run 8025 Trigger/ROC Snapshots
  • 337282 10/29/10 06:05 lselvy Target Update
  • 337280 10/29/10 05:11 flay Run 8024 Trigger/ROC Snapshots
  • 337269 10/29/10 03:31 flay Run 8021 Trigger/ROC Snapshots
  • 337263 10/29/10 02:48 jroche propagated pion rejector map to online display golden run
  • 337260 10/29/10 02:22 flay Run 8020 Trigger/ROC Snapshots
  • 337259 10/29/10 02:05 jroche Calorimeter DC current
  • 337256 10/29/10 00:42 jroche Helicity flip rate restored to 943 Hz
  • 337251 10/29/10 00:13 flay Run 8017 Trigger/ROC Snapshots
  • 337249 10/29/10 00:00 M Friend Helicity board set to 120Hz, Compton DAQ set to 1kHz
  • Thursday

  • 337248 10/28/10 23:59 efuchey 10/28/10 Swing shift summary
  • 337247 10/28/10 23:38 cshloka Pion rejector mapping fixed
  • 337246 10/28/10 23:37 LeRose Target Screens
  • 337243 10/28/10 23:08 itard entries end of run
  • 337242 10/28/10 22:52 efuchey Problem with Compton online analyzer
  • 337239 10/28/10 22:08 itard entries end of run
  • 337237 10/28/10 21:06 LeRose Target Screens
  • 337236 10/28/10 21:06 itard entries end of run
  • 337234 10/28/10 20:40 Carlos Deadtime in buffered mode
  • 337231 10/28/10 20:06 itard dead time study 4.5 uA
  • 337223 10/28/10 19:50 itard dead time study 3uA
  • 337218 10/28/10 19:21 itard dead time study 2 uA
  • 337217 10/28/10 19:12 efuchey current change
  • 337208 10/28/10 18:53 LeRose Target Screens
  • 337206 10/28/10 18:53 itard dead time study
  • 337203 10/28/10 18:38 efuchey Cavity scan attempt
  • 337201 10/28/10 18:27 itard logs of end of run
  • 337199 10/28/10 17:53 Carlos Dead time study with buffering
  • 337196 10/28/10 17:41 efuchey Compton Rates for 10-28-10
  • 337195 10/28/10 17:29 Carlos Pedestals fluctuations
  • 337192 10/28/10 17:22 efuchey Hall A tools screenshot
  • 337191 10/28/10 17:20 efuchey Calo base currents
  • 337188 10/28/10 17:10 LeRose Target to LH2
  • 337186 10/28/10 16:56 efuchey Harp scan
  • 337185 10/28/10 16:51 efuchey Beam is back
  • 337184 10/28/10 16:34 efuchey Hall in beam permit
  • 337182 10/28/10 16:30 A. Camsonne DAQ Sync work
  • 337170 10/28/10 15:01 A. Camsonne ARS firmware version E12 loaded
  • 337168 10/28/10 15:00 Carlos Pedestal fluctuations
  • 337158 10/28/10 14:17 Fansler Moeller
  • 337156 10/28/10 13:46 cshloka target screens snapshot
  • 337155 10/28/10 13:19 Jack Pion Rejector cabling
  • 337149 10/28/10 11:50 efuchey QWP scan for 10-28-10
  • 337147 10/28/10 11:15 O.Glamazdin Beam polarization sign
  • 337146 10/28/10 10:57 Fansler/Ewin RQ2 04/16 Valves
  • 337140 10/28/10 10:43 A. Camsonne Fastbus ADC vs V792 ADC in DVCS crate
  • 337139 10/28/10 10:19 jroche testing the end of run scripts
  • 337135 10/28/10 10:07 A. Camsonne /adaql4/work1 full removing PREX rootfiles
  • 337131 10/28/10 09:40 cshloka target status and L2 cernox fluctuations
  • 337127 10/28/10 07:25 lselvy Target Move
  • 337126 10/28/10 07:23 Hisham Controlled access
  • 337125 10/28/10 07:15 Hisham Owl shift summary
  • 337124 10/28/10 07:14 flay Run 7987 is a Junk Run
  • 337123 10/28/10 07:10 flay Run 7986 is a Junk Run
  • 337122 10/28/10 07:07 flay Run 7986 is a Junk Run
  • 337118 10/28/10 07:02 flay Run 7985 Trigger/ROC Snapshots
  • 337117 10/28/10 06:05 lselvy Target Update
  • 337114 10/28/10 06:01 flay Run 7984 Trigger/ROC Snapshot
  • 337113 10/28/10 05:03 flay Run 7984 Trigger/ROC Snapshots
  • 337110 10/28/10 04:01 flay Run 7983 Trigger/ROC Snapshots
  • 337102 10/28/10 03:02 flay Run 7982 Trigger/ROC Snapshots
  • 337093 10/28/10 02:03 flay Run 7981 Trigger/ROC Snapshots
  • 337091 10/28/10 01:52 hisham Hall A General Tools
  • 337090 10/28/10 01:50 hisham PbF2 Calo ADC Readback
  • 337089 10/28/10 01:46 Hisham dvcstlab1.2006 Alarm
  • 337088 10/28/10 01:02 flay Run 7980 Trigger/ROC Snapshots
  • Wednesday

  • 337083 10/27/10 23:59 A. Camsonne Readout of ROC18 and trigger library updated
  • 337082 10/27/10 23:59 Yi Qiang Swing Shift Summary
  • 337080 10/27/10 23:40 mcanan swing_target_screens
  • 337077 10/27/10 23:34 Punjabi roc screens
  • 337076 10/27/10 23:32 Punjabi scaler, RC screens
  • 337075 10/27/10 23:16 Yi Qiang Calorimeter current
  • 337074 10/27/10 23:16 efuchey Rates became too high
  • 337072 10/27/10 22:39 Yi Qiang Problem solved, back to production
  • 337071 10/27/10 22:24 Yi Qiang Beam will be off for a while
  • 337070 10/27/10 22:19 Yi Qiang Beam energy EPICS value missing at the end of run 7975
  • 337066 10/27/10 22:00 A. Camsonne Left HRS efficiency run Run 7975
  • 337063 10/27/10 21:55 Punjabi roc screens
  • 337062 10/27/10 21:54 Punjabi scaler and RC screens
  • 337061 10/27/10 21:19 efuchey Compton CODA crash
  • 337058 10/27/10 21:11 Punjabi roc screens
  • 337057 10/27/10 21:11 Punjabi RC, Scaler screens
  • 337054 10/27/10 20:26 Punjabi Roc screens
  • 337053 10/27/10 20:25 Punjabi rc, scaler screens
  • 337043 10/27/10 19:18 Punjabi roc screens
  • 337042 10/27/10 19:16 Punjabi roc 14 and roc 17 screens
  • 337041 10/27/10 19:14 Punjabi RC and scaler screens
  • 337040 10/27/10 19:01 efuchey Compton rates for 10-27-10
  • 337033 10/27/10 18:27 mcanan Target Motion
  • 337032 10/27/10 18:00 mcanan Target Motion
  • 337031 10/27/10 17:16 Yi Qiang No beam for a while due to magnet trips
  • 337030 10/27/10 17:00 A. Camsonne ARS firmware version E10 downloaded in 12 ARS boards
  • 337028 10/27/10 16:44 Yi Qiang Beam back to the Hall, continue production
  • 337027 10/27/10 16:38 mcanan Target Motion
  • 337025 10/27/10 15:59 Jixie Moller has been refilled
  • 337024 10/27/10 15:57 Saha Shift Summary
  • 337023 10/27/10 15:55 O.Glamazdin Moller results
  • 337022 10/27/10 15:45 Lamiaa Target's Monitoring tools
  • 337021 10/27/10 15:42 Missing proton mass
  • 337019 10/27/10 13:33 A. Camsonne Run7967 Junk run
  • 337014 10/27/10 13:06 A. Camsonne Loaded tsUtil.o on TS11 crate
  • 337010 10/27/10 12:39 A. Camsonne Run 7964 run in interrupt mode
  • 337006 10/27/10 12:22 A. Camsonne Empty FIFO messages Run 7692 with buffering
  • 337004 10/27/10 12:15 A. Camsonne ARS Firmware E11 loaded for buffering tests - HV on cosmics paddles on
  • 337001 10/27/10 11:51 A. Camsonne DVCS calorimeter HV off for Moller measurement
  • 336990 10/27/10 10:39 O.Glamazdin Moller magnet is rumping up
  • 336989 10/27/10 10:14 efuchey shift check list modified
  • 336988 10/27/10 09:58 gaskell Run plan for DAY and SWING shift
  • 336987 10/27/10 09:50 efuchey QWP scan for 10-27-10
  • 336986 10/27/10 09:26 O.Glamazdin Controlled acces for the Moller solenoid reboot
  • 336985 10/27/10 09:17 Lamiaa Strange behavior of Hal2_Cernox_2_R
  • 336984 10/27/10 08:44 Lamiaa Screenshot of target's charts
  • 336983 10/27/10 08:34 Lamiaa Move target to Empty position
  • 336981 10/27/10 08:11 jroche Owl shift summary
  • 336977 10/27/10 07:54 flay Run 7959 Trigger/ROC Snapshot
  • 336975 10/27/10 07:11 flay Run 7958 is a Junk Run
  • 336972 10/27/10 07:09 flay Run 7957 is a Junk Run
  • 336968 10/27/10 06:18 flay Run 7956 Trigger/ROC Snapshot
  • 336965 10/27/10 05:51 Carlos Pedestal issues with new ARS firmware (E11)
  • 336964 10/27/10 05:47 jroche Increased compton HV to 1675V
  • 336963 10/27/10 05:46 jroche It's freezing in here !
  • 336959 10/27/10 05:04 A. Camsonne ARS firmware version E10 downloaded in 12 ARS boards
  • 336957 10/27/10 04:59 flay Some CALO Blocks Not Working
  • 336956 10/27/10 04:48 A. Camsonne DVCS sync check work
  • 336942 10/27/10 03:46 jroche Compton rates look reasonnable
  • 336941 10/27/10 03:46 itard target graph
  • 336939 10/27/10 03:28 jroche Calo anode currents
  • 336937 10/27/10 03:23 jroche general Tools at the beginning of the data taking
  • 336936 10/27/10 03:20 jroche harps looks good
  • 336934 10/27/10 03:02 jroche data taking to resume shortly
  • 336897 10/27/10 00:01 GOLGE Swing Shift Summary
  • Tuesday

  • 336880 10/26/10 22:07 Golge Controlled access
  • 336875 10/26/10 20:33 A. Camsonne Networked FPGA programmer replugged
  • 336865 10/26/10 19:20 gaskell Run plan for Tuesday Swing --> Wednesday Day
  • 336852 10/26/10 18:53 Magali news trigger fpga ctrl firmware
  • 336850 10/26/10 18:46 A. Camsonne ROC18 power cycled
  • 336846 10/26/10 18:36 efuchey Compton rates for 10-26-10
  • 336838 10/26/10 18:14 jroche BCM calibration: some result.
  • 336836 10/26/10 18:01 mcanan Target Motion
  • 336835 10/26/10 17:44 mcanan Target Motion
  • 336834 10/26/10 17:42 A. Camsonne DVCS calorimeter cosmics paddle off
  • 336833 10/26/10 17:15 A. Camsonne Target operator refresher by JP on DVCS wiki
  • 336832 10/26/10 16:24 A. Camsonne APC src2 channel assignement
  • 336831 10/26/10 16:19 A. Camsonne iochacs serial port on dvcstlab1 2008
  • 336829 10/26/10 16:03 Holmstrom Day Shift Summary
  • 336828 10/26/10 16:02 A. Camsonne Slow control APC PDU switched to src2 hostname
  • 336827 10/26/10 15:48 A. Camsonne Calorimeter cosmics paddles HV on for buffering test
  • 336825 10/26/10 15:42 A. Camsonne ARS Firware version E11
  • 336819 10/26/10 14:22 Jixie Moller magnet has been filled
  • 336818 10/26/10 13:19 Carlos HRS moved to new angle (31.82 deg: kin3 low)
  • 336817 10/26/10 12:27 hyde S0 Efficiency
  • 336810 10/26/10 11:27 efuchey Compton beamline
  • 336809 10/26/10 11:12 Fansler/Ewin Left Dipole
  • 336801 10/26/10 10:32 gaskell Status and plan for the day
  • 336800 10/26/10 10:29 gaskell HRS work in progress
  • 336799 10/26/10 10:07 Holmstrom Spectrometer Off
  • 336798 10/26/10 09:14 gomez Re: Problem cycling quadrupole current
  • 336796 10/26/10 08:19 Hisham Run 7895: ROC 18 and 14
  • 336795 10/26/10 08:17 Hisham Run 7895 scalers, RunControl gui
  • 336794 10/26/10 08:11 Hisham Hall in controlled access
  • 336793 10/26/10 08:09 Hisham Owl shift summary
  • 336790 10/26/10 07:46 cshloka target status
  • 336789 10/26/10 07:34 flay Run 7894 is Junk
  • 336785 10/26/10 07:02 flay Run 7893 Trigger/ROC Snapshots
  • 336783 10/26/10 06:15 flay Run 7892 Physics Quantities
  • 336782 10/26/10 06:12 flay Run 7892 Calo TDCs
  • 336781 10/26/10 06:09 flay Run 7892 LHRS Gas Cherenkov
  • 336780 10/26/10 06:07 flay Run 7892 S1 TDCs
  • 336779 10/26/10 05:57 flay Run 7892 Trigger/ROC Snapshots
  • 336778 10/26/10 05:13 cshloka target status and L2 cernox
  • 336776 10/26/10 04:50 flay Run 7891 is Junk
  • 336775 10/26/10 04:43 Hisham Hall A General Tools snapshot.
  • 336774 10/26/10 04:37 flay Run 7890 LHRS GC
  • 336772 10/26/10 04:13 flay Run 7890 Trigger/Snapshots
  • 336766 10/26/10 03:35 flay Run 7889 Exhibits Same Behavior as Run 7888
  • 336764 10/26/10 03:20 Hisham Hall A General Tools snapshot
  • 336763 10/26/10 03:14 flay Run 7889: GC Sum
  • 336761 10/26/10 02:59 flay Run 7889 Trigger/ROC Snapshots
  • 336755 10/26/10 02:34 Carlos Calorimeter HV update
  • 336754 10/26/10 02:29 Hisham Problems in ycling quadrupole currents
  • 336752 10/26/10 02:09 cshloka target status
  • 336751 10/26/10 01:53 flay Important: Database Updated
  • 336750 10/26/10 01:17 Hisham LHRS moved to 37.86 deg
  • 336749 10/26/10 01:11 flay Run 7888 Physics Quantities
  • 336748 10/26/10 01:08 flay Run 7888 Calorimeter TDCs
  • 336747 10/26/10 01:01 flay Run 7888
  • 336745 10/26/10 00:14 flay Run 7888 Trigger/ROC Snapshot
  • 336744 10/26/10 00:05 Min Huang target charts screen
  • 336742 10/26/10 00:03 BoZhao Swing shift summary
  • Monday

  • 336740 10/25/10 23:25 Margaziotis Run 7887 ROC18 screen
  • 336739 10/25/10 23:24 Margaziotis Run 7887 ROC14 and ROC17 screens
  • 336738 10/25/10 23:23 Margaziotis Run 7887 CODA and scaler screens
  • 336736 10/25/10 22:51 Min Huang target moved to loop2, 15cm
  • 336735 10/25/10 22:39 BoZhao 4A and 4B Harp scan
  • 336731 10/25/10 21:07 Magali new firmware version
  • 336730 10/25/10 21:06 efuchey Compton tuning
  • 336726 10/25/10 19:52 BoZhao Stop taking data for compton tuning
  • 336720 10/25/10 19:47 Min Huang move to empty target
  • 336719 10/25/10 19:46 efuchey Compton rates at another volage
  • 336718 10/25/10 19:40 Margaziotis Run 7884 ROC18 screen
  • 336717 10/25/10 19:39 Margaziotis Run 7884 ROC14 and ROC17
  • 336716 10/25/10 19:38 Margaziotis Run 7884 CODA and scaler screens
  • 336714 10/25/10 19:25 efuchey Compton rates 10-25-10
  • 336712 10/25/10 19:00 Margaziotis Run 7883 ROC18 screen
  • 336711 10/25/10 19:00 Margaziotis Run 7883 ROC14 and ROC17 screens
  • 336710 10/25/10 18:59 Margaziotis Run 7883 CODA and scalers screen
  • 336707 10/25/10 18:16 Margaziotis Run 7882 ROC 18 screen
  • 336706 10/25/10 18:15 Margaziotis Run 7882 ROC14 and ROC17 screens
  • 336705 10/25/10 18:14 Margaziotis Run 7882 CODA and scaler screens
  • 336703 10/25/10 17:31 BoZhao Hall A general tools screen
  • 336702 10/25/10 17:27 Jack (Techs) DVCS grounding strap is too short
  • 336700 10/25/10 16:24 A. Camsonne Calorimeter ADCs channels in end of run time outs
  • 336699 10/25/10 16:21 gaskell Run plan for Monday Day and Swing
  • 336697 10/25/10 16:08 JPChen Day shit summary
  • 336694 10/25/10 15:30 O.Glamazdin Run 7880: ROC1
  • 336693 10/25/10 15:30 O.Glamazdin Run 7880: ROC14 and ROC17
  • 336692 10/25/10 15:29 O.Glamazdin Run 7880: CODA and Scalers
  • 336691 10/25/10 15:16 R. Michaels added roc14,17 scalers to datastream
  • 336689 10/25/10 15:07 R. Michaels xscaler updated, can display ROC14, 17
  • 336688 10/25/10 14:59 jpchen HRSL to elastic 2b
  • 336686 10/25/10 14:34 O.Glamazdin Run 7879: ROC18
  • 336685 10/25/10 14:33 O.Glamazdin Run 7879: ROC14 and ROC17
  • 336684 10/25/10 14:32 O.Glamazdin Run 7879: CODA and Scalers
  • 336683 10/25/10 14:11 O.Glamazdin Calorimeter HV
  • 336680 10/25/10 13:38 Hisham s1 and s2 efficiency
  • 336679 10/25/10 13:32 O.Glamazdin Run 7878: ROC18
  • 336678 10/25/10 13:32 O.Glamazdin Run 7878: ROC14, ROC17
  • 336677 10/25/10 13:30 O.Glamazdin Run 7878 Coda and Scaler ratesd
  • 336676 10/25/10 13:06 Hisham s2 efficiency with the right figure
  • 336675 10/25/10 13:01 R. Michaels datamon updated again (use T10)
  • 336674 10/25/10 12:51 Hisham s2 efficiency
  • 336672 10/25/10 12:44 jpchen Harp 1H04A
  • 336671 10/25/10 12:43 jpchen Harp 1H04B
  • 336670 10/25/10 12:31 Carlos HV update
  • 336669 10/25/10 12:20 gaskell Unser monitor temp control bad?
  • 336668 10/25/10 12:16 jpchen request beam permit
  • 336667 10/25/10 12:14 A. Camsonne Scope traces block checks
  • 336666 10/25/10 12:13 Magali trigger firmware
  • 336665 10/25/10 11:53 A. Camsonne Scope traces block checks
  • 336664 10/25/10 11:52 jpchen Back to controlled access
  • 336663 10/25/10 11:24 hyde S1 HV adjustment
  • 336662 10/25/10 11:18 Jack Pion Rejector cabling
  • 336661 10/25/10 11:06 jpchen controlled access over, back to beam permit
  • 336660 10/25/10 11:05 jpchen calorimeter and HRSL moved and set for ealstic 2-a
  • 336657 10/25/10 10:16 A. Camsonne Left HRS positive polarity and set to 2.245 GeV/c
  • 336656 10/25/10 09:51 A. Camsonne Calorimeter cosmics HV from 600 V to 700 V
  • 336655 10/25/10 09:42 A. Camsonne Run 7876 cosmics run with buffering
  • 336653 10/25/10 09:40 O.Glamazdin Moller magnet has been filled
  • 336650 10/25/10 09:13 jpchen In controlled access
  • 336649 10/25/10 08:45 jpchen Controlled access to change to Elastic 2 (a) setting
  • 336648 10/25/10 08:34 R. Michaels datamon updated for DVCS
  • 336647 10/25/10 08:29 jroche turning on and off various pion rejector HV
  • 336645 10/25/10 08:06 jroche owl shift summary
  • 336644 10/25/10 07:50 Markowitz moving to empty target, LHRS ramping down
  • 336643 10/25/10 07:41 jroche pion rejector mystery solved?
  • 336642 10/25/10 07:33 lamiaa Rocs[14,17&18] screenshot of a run 7874
  • 336640 10/25/10 07:15 Lamiaa Recapitulation of CODA Screenshot of a run 7873
  • 336639 10/25/10 07:13 Lamiaa Recapitulation of Screenshot of Rocs[14,17&18] of a run 7873
  • 336637 10/25/10 06:49 Lamiaa CODA screenshot of a run 7873
  • 336636 10/25/10 06:44 Lamiaa Screenshot of Rocs[14,17&18] of a run 7873
  • 336634 10/25/10 06:22 Lamiaa Late halog entry of the end of run #7867
  • 336632 10/25/10 06:07 jroche all HV back on for the pion rejector
  • 336629 10/25/10 06:02 Lamiaa Screenshot of CODA & Scalers of a run 7872
  • 336628 10/25/10 06:01 Lamiaa Screenshot of Rocs[14,17&18] of a run 7872
  • 336626 10/25/10 05:21 jroche only prl2 ch 12 and 13 off
  • 336624 10/25/10 05:16 Lamiaa Screenshot of CODA & Scalers of a run 7871
  • 336623 10/25/10 05:15 Lamiaa Screenshot of Rocs[14,17&18] of a run 7871
  • 336620 10/25/10 04:29 Lamiaa Screenshot of CODA & Scalers of a run 7870
  • 336619 10/25/10 04:29 Lamiaa Screenshot of Rocs[14,17&18] of a run 7870
  • 336618 10/25/10 03:53 Lamiaa Screenshot of CODA & Scalers of a run 7869
  • 336617 10/25/10 03:51 lamiaa Screenshot of Rocs[14&17] of a run 7869
  • 336614 10/25/10 03:41 jroche more pion rejector off
  • 336611 10/25/10 03:36 Lamiaa Screenshot of CODA & Scalers of a run 7868
  • 336610 10/25/10 03:35 Lamiaa Screenshot of Rocs[14&17] of a run 7868
  • 336609 10/25/10 03:20 jroche prl signals during run 7867
  • 336608 10/25/10 03:05 jroche some prl turned off (the screen)
  • 336607 10/25/10 03:04 jroche other Pion rejector turned off.
  • 336605 10/25/10 02:50 Lamiaa Screenshot of DOCA & Scalers of a run 7867
  • 336604 10/25/10 02:50 Lamiaa Screenshot of Rocs[14&17] of a run 7867
  • 336603 10/25/10 02:16 jroche pion rejector: some pion rejector off during run 7867
  • 336601 10/25/10 02:07 jroche change pedestal for S2R15
  • 336598 10/25/10 02:02 Lamiaa Screenshot of CODA & Scalers of a run 7866
  • 336597 10/25/10 02:01 Lamiaa Screenshot of Rocs[14&17] of a run 7866
  • 336596 10/25/10 01:52 Markowitz target screens
  • 336595 10/25/10 01:35 Lamiaa S0 and S2 distributions of a run 7865
  • 336593 10/25/10 01:14 Carlos Calorimeter work
  • 336591 10/25/10 00:50 Markowitz target to empty and back to LH2
  • Sunday

  • 336589 10/24/10 23:56 Margaziotis Run 7864 ROC18 screenshot
  • 336588 10/24/10 23:56 Margaziotis Run 7864 ROC14 and ROC17 screenshots
  • 336587 10/24/10 23:54 Margaziotis Run 7854 CODA and Scaler screens
  • 336585 10/24/10 23:54 Ramesh Subed Shift Summary
  • 336584 10/24/10 23:53 Margaziotis Run 7864 HRS spectra need attention
  • 336582 10/24/10 23:52 Ramesh Asked for short escorted access
  • 336581 10/24/10 23:44 Margaziotis Run 7864 Calorimeter ADC and TDC spectra empty
  • 336580 10/24/10 23:38 Margaziotis Run 7864 Calorimeter ADC 84 spectrum empty
  • 336579 10/24/10 23:22 Carlos T10 fixed
  • 336574 10/24/10 23:08 BoZhao Target charts screen
  • 336573 10/24/10 23:03 BoZhao Target to LH2
  • 336572 10/24/10 22:44 BoZhao Target to Empty
  • 336571 10/24/10 22:39 Ramesh Hall in beam permit
  • 336554 10/24/10 20:08 Jixie Moller magnet filling was unusual slow today
  • 336553 10/24/10 20:05 A. Camsonne vxserver loaded on iochacs and dvcstlab6
  • 336552 10/24/10 20:05 jroche, sega pion rejector cabling check
  • 336549 10/24/10 20:00 Hisham BNC synch cable
  • 336548 10/24/10 19:54 J. Zhang Moller magnet has been filled
  • 336543 10/24/10 19:29 Carlos Calo HV ON
  • 336539 10/24/10 19:11 gaskell Run plan for Sunday swing to Monday Day
  • 336537 10/24/10 18:08 Carlos Calorimeter HV off
  • 336536 10/24/10 18:01 jroche more on the pion rejector
  • 336534 10/24/10 17:40 Ramesh Asked for controlled access
  • 336533 10/24/10 17:37 Margaziotis Run 7846 ROC18
  • 336532 10/24/10 17:37 Margaziotis Run 7846 ROC14 and ROC17 screen
  • 336531 10/24/10 17:35 Margaziotis Run 7846 ROC14 and ROC17
  • 336530 10/24/10 17:34 Margaziotis Run 7846 CODA and scalers screenshot
  • 336529 10/24/10 17:26 BoZhao target charts screen
  • 336528 10/24/10 17:13 Carlos 2-photon invariant mass
  • 336527 10/24/10 17:11 jroche one more fix to the end of run epics script (better formatting)
  • 336526 10/24/10 17:11 jroche one more fix to the end of run epics script
  • 336523 10/24/10 17:02 Margaziotis Run 7845 ROC14 and ROC17 screenshots
  • 336522 10/24/10 17:02 Margaziotis Run 7845 CODA and scalers screenshot
  • 336519 10/24/10 16:17 Carlos Calo hit map
  • 336518 10/24/10 16:12 Margaziotis run 7844 ROC14 and ROC17
  • 336516 10/24/10 16:11 Margaziotis Run 7844 CODA and scaler screenshots
  • 336515 10/24/10 16:05 Lindgren Shift summary
  • 336514 10/24/10 16:03 BoZhao Target charts screen
  • 336513 10/24/10 15:36 Lamiaa LHRS Cerenkov signal on PMT2 and sum back to normal on previous run 7843!
  • 336512 10/24/10 15:36 Carlos Run 7844 with DVCSTopBot configuration
  • 336511 10/24/10 15:30 Lindgren current strip chart
  • 336510 10/24/10 15:29 Lindgren HV DVCS
  • 336508 10/24/10 15:28 Lindgren DVCS DC Levels
  • 336506 10/24/10 15:26 Lindgren Hall A Tools
  • 336505 10/24/10 15:26 Lamiaa Screenshot of CODA & Scalers of a run 7843
  • 336504 10/24/10 15:25 Lamiaa Screenshot of Rocs[14,17&18] of a run 7843
  • 336503 10/24/10 15:11 jroche Modified epics run start and end run to make them faster
  • 336501 10/24/10 15:00 Lamiaa Cerenkov data deviation on PMT2 and sum noticed on runs 7841 & 7842
  • 336500 10/24/10 14:57 gaskell Controlled access at 6:15 pm
  • 336494 10/24/10 14:40 Lamiaa Screenshot of CODA & Scalers of a run 7842
  • 336493 10/24/10 14:37 Lamiaa Screenshot of Rocs[14,17&18] of a run 7842
  • 336492 10/24/10 14:29 M trigger informations
  • 336491 10/24/10 14:24 Magali firmware of the calorimeter trigger
  • 336488 10/24/10 13:51 Lamiaa Screenshot of CODA & Scalers of a run 7841
  • 336487 10/24/10 13:50 Lamiaa Screenshot of Rocs[14,17&18] of a run 7841
  • 336484 10/24/10 13:07 Lamiaa Screenshot of CODA & Scalers of a run 7840
  • 336483 10/24/10 13:06 Lamiaa Screenshot of Rocs[14,17&18] of a run 7840
  • 336482 10/24/10 12:51 Carlos Dead time reading
  • 336481 10/24/10 12:33 Lindgren Lower beam current
  • 336478 10/24/10 12:26 Lindgren Half shift summary
  • 336477 10/24/10 12:20 Lamiaa Screenshot of CODA & Scalers of a run 7839
  • 336476 10/24/10 12:19 Lamiaa Screenshot of Rocs[14,17&18] of a run 7839
  • 336475 10/24/10 12:16 Lamiaa Missing entry of the "End of run 7838"!
  • 336474 10/24/10 12:09 Lamiaa Real Start time is 11:34AM instead of the shown posting time 11:51AM
  • 336473 10/24/10 12:05 Carlos Start/End of run scripts
  • 336470 10/24/10 11:33 Lamiaa Screenshot of CODA and Scalers of a run 7838
  • 336469 10/24/10 11:32 Lamiaa Screenshot of Rocs[14,17&18] of a run 7838
  • 336468 10/24/10 11:27 Lamiaa Halog timing issue!
  • 336467 10/24/10 11:12 Carlos ROC screenshots
  • 336464 10/24/10 10:36 Lamiaa Screenshot of CODA and Scalers of a run 7837
  • 336463 10/24/10 10:35 Lamiaa Screenshot of Rocs[14,17&18] of a run 7837
  • 336462 10/24/10 10:11 gaskell Hall C energy modulation
  • 336459 10/24/10 09:51 lamiaa Screenshot of Rocs[14,17&18] by the end of a run 7836
  • 336458 10/24/10 09:49 lamiaa Screenshot of CODA and Scalers of a run 7836
  • 336457 10/24/10 09:23 Lamiaa Screenshot of Roc17, Roc18 & Roc 14 of run 7836
  • 336453 10/24/10 08:39 Lamiaa Screenshot of CODA and Scaler of a run 7833
  • 336452 10/24/10 08:35 Lamiaa Screenshot of Roc17, Roc18 & Roc14 of a run 7833
  • 336449 10/24/10 08:16 Lindgren Starting Day shift
  • 336448 10/24/10 08:01 efuchey 10/24/10 Owl shift summary
  • 336444 10/24/10 07:08 D.Wang rocs
  • 336443 10/24/10 07:07 D.Wang Coda and scalers
  • 336439 10/24/10 05:23 D.Wang rocs
  • 336438 10/24/10 05:22 D.Wang coda and scalers
  • 336437 10/24/10 04:54 efuchey DVCS HV alarm
  • 336433 10/24/10 04:32 D.Wang rocs
  • 336432 10/24/10 04:31 D.Wang Coda and scalers
  • 336429 10/24/10 03:42 D.Wang rocs
  • 336428 10/24/10 03:42 D.Wang coda and scalers
  • 336425 10/24/10 02:56 D.Wang rocs
  • 336424 10/24/10 02:56 D.Wang coda and scalers
  • 336422 10/24/10 02:20 efuchey HV trip: complementary information.
  • 336420 10/24/10 02:16 M. Magne Current logic module version
  • 336412 10/24/10 01:54 efuchey HV trip
  • 336408 10/24/10 01:30 A. Camsonne Running config from now on TrigDVCS
  • 336406 10/24/10 01:20 A. Camsonne trigROC18.crl modified to read 127 words TrigDVCS modified - Run 7821
  • 336404 10/24/10 01:07 A. Camsonne Added dvcs3 flag to DVCS3 config - starting run 7819
  • 336403 10/24/10 01:01 A. Camsonne Run 7818 Run with ROC18 trigger and Top and Bottom crate
  • 336400 10/24/10 00:52 D.Wang rocs
  • 336399 10/24/10 00:52 D.Wang coda and scalers
  • 336398 10/24/10 00:32 efuchey PbF2 base currents
  • 336397 10/24/10 00:31 efuchey Hall A tools screenshot
  • 336394 10/24/10 00:01 Margaziotis Run 7815 ROC14 and ROC17 screenshots
  • 336392 10/24/10 00:00 Margaziotis Run 7815 DAQ screenshot
  • 336391 10/24/10 00:00 GOLGE shift summary for swing
  • Saturday

  • 336390 10/23/10 23:59 Margaziotis Run 7815 scalers screenshot
  • 336389 10/23/10 23:45 Carlos T10 scaler
  • 336388 10/23/10 23:43 Margaziotis HRS spectra need checking
  • 336386 10/23/10 23:16 Maargaziotis Run 7814 ROC14 and ROC17 screenshots
  • 336385 10/23/10 23:15 Margaziotis Run 7814 DAQ screenshot
  • 336384 10/23/10 23:14 Margaziotis Run 7814 scalers screenshot
  • 336379 10/23/10 21:55 Margaziotis Run 7812 ROC14 and ROC17 screenshots
  • 336378 10/23/10 21:54 Margaziotis Run 7812 runcontrol screenshot
  • 336377 10/23/10 21:53 Margaziotis Run 7812 scalers screenshot
  • 336374 10/23/10 21:13 Carlos HV change: block 119
  • 336372 10/23/10 21:01 Margaziotis Restarted Coda to fix bug
  • 336371 10/23/10 20:51 Margaziotis Run 7809 Rock14 and Rock17 screenshots
  • 336370 10/23/10 20:49 Margaziotis Run 7809 runcontrol screenshot
  • 336369 10/23/10 20:47 Margaziotis Run 7809 scalers screenshot
  • 336368 10/23/10 20:37 Margaziotis Calo Amp/TDC 119 spectra run 7809
  • 336366 10/23/10 20:02 ZAFAR AHMED Target is LH2 Loop 2 15 cm
  • 336365 10/23/10 19:56 R. Michaels re: work disk cleanup
  • 336362 10/23/10 19:29 Carlos Block 119 bloken, fixed now
  • 336361 10/23/10 19:23 A. Camsonne HV cosmics off - reloaded HV from Oct 19
  • 336352 10/23/10 18:47 A. Camsonne Changed sync event frequency to 10
  • 336350 10/23/10 18:24 A. Camsonne Nothing on BNC sync signal
  • 336348 10/23/10 18:23 A. Camsonne Buffering making empty FIFO at 60 Hz of cosmics
  • 336347 10/23/10 18:16 A. Camsonne crl for ROC17 and ROC14 modified
  • 336334 10/23/10 16:54 A. Camsonne ROC18 hardware ROC ID switched from 5 to 1
  • 336331 10/23/10 16:19 J. Zhang Moller magnet has been filled
  • 336330 10/23/10 16:04 Lindgren Shift summary
  • 336325 10/23/10 15:38 A. Camsonne Test with validation - need to check timing
  • 336320 10/23/10 15:21 A. Camsonne Run 7787 test with buffering
  • 336319 10/23/10 14:58 A. Camsonne ROC1417Buff modified for buffering
  • 336316 10/23/10 14:52 chench scalers and DAQ of run7786
  • 336315 10/23/10 14:51 chench Screenshot of ROC14 and ROC17 for run7786
  • 336312 10/23/10 14:03 chench Screenshot of ROC14 and ROC17 for run7785
  • 336311 10/23/10 14:02 chench Scalers of run7785
  • 336310 10/23/10 14:02 chench Screenshot of DAQ for run#7785
  • 336309 10/23/10 14:00 Lindgren DVCS Calo
  • 336308 10/23/10 13:59 lindgren Lecroy HV
  • 336307 10/23/10 13:57 Lindgren beam chart
  • 336306 10/23/10 13:54 Lindgren Hall A Tools
  • 336305 10/23/10 13:53 Lindgren DVCS ADC Readback
  • 336304 10/23/10 13:52 Ole Hansen Several work1 disks are full/nearly full
  • 336303 10/23/10 13:40 chench Missing channels of pioin rejectors1 ADC
  • 336302 10/23/10 13:36 markowit target screens
  • 336301 10/23/10 13:36 Ole Hansen Don't fill adaq home directory quota!
  • 336298 10/23/10 13:10 Margaziotis Screenshot DAQ Run 7784
  • 336297 10/23/10 13:09 Margaziotis ROCK14 &ROCK17 run 7784
  • 336296 10/23/10 13:08 Margaziotis Screenshot for scalers run 7784
  • 336293 10/23/10 12:13 chench Screenshot for DAQ of run7783
  • 336292 10/23/10 12:12 chench Screenshot for ROC14 and ROC17 of run7783
  • 336291 10/23/10 12:11 chench Scalers for run7783
  • 336288 10/23/10 11:18 chench screen shot of ROC14 and ROC17 for run7782
  • 336287 10/23/10 11:17 chench Screen shot of DAQ for run7782
  • 336286 10/23/10 11:16 chench Scalers of s1, s2 for run7782
  • 336285 10/23/10 11:06 A. Camsonne Can you also take the scalers norm tab ?
  • 336283 10/23/10 10:36 gaskell Access scheduled for 3 pm
  • 336282 10/23/10 10:25 chench Screenshot of DAQ for run7781
  • 336281 10/23/10 10:24 chench Scalers for run7781
  • 336280 10/23/10 10:23 chench Roc14 and Roc17 screenshot for run7781
  • 336279 10/23/10 10:01 A. Camsonne Issues with power cycling the crates
  • 336278 10/23/10 10:01 A. Camsonne Issues with power cycling the crates
  • 336277 10/23/10 09:43 Lindgren Coda failure
  • 336275 10/23/10 09:12 chench Error from DAQ during run7780
  • 336272 10/23/10 08:50 chench Roc14 and Roc17 screen shot at the end of run7779
  • 336271 10/23/10 08:14 LIndgren Statring Day Shift
  • 336267 10/23/10 07:59 efuchey Owl 10-23-10 shift summary
  • 336266 10/23/10 07:49 D.Wang rocs
  • 336265 10/23/10 07:49 D.Wang coda and scalers
  • 336262 10/23/10 06:56 D.Wang rocs
  • 336261 10/23/10 06:56 D.Wang coda and scalers
  • 336258 10/23/10 06:06 D.Wang rocs
  • 336257 10/23/10 06:05 D.Wang coda and scalers
  • 336256 10/23/10 05:44 efuchey PbF2 base currents
  • 336254 10/23/10 05:21 efuchey Target screens
  • 336252 10/23/10 05:20 efuchey Beam back
  • 336251 10/23/10 05:10 D.Wang rocs
  • 336250 10/23/10 05:10 D.Wang Coda and scalers
  • 336247 10/23/10 03:43 efuchey Beqam down again
  • 336246 10/23/10 03:26 efuchey Hall C feedback test done
  • 336243 10/23/10 03:09 efuchey Beam is back
  • 336239 10/23/10 02:47 efuchey Beam down for more than 15 minutes
  • 336238 10/23/10 02:23 D.Wang rocs
  • 336237 10/23/10 02:23 D.Wang coda and scalers
  • 336236 10/23/10 01:54 efuchey Hall A tools screenshot
  • 336233 10/23/10 01:32 D.Wang rocs
  • 336232 10/23/10 01:32 D.Wang coda and scaler
  • 336231 10/23/10 01:16 efuchey Compton rates for tonight
  • 336228 10/23/10 00:46 D.Wang rocs
  • 336227 10/23/10 00:45 D.Wang runcontrol and scalers
  • 336225 10/23/10 00:03 ZAFAR AHMED Friday 10-22-2010 Swing Shift Summary
  • Friday

  • 336222 10/22/10 23:42 Nuruzzaman Target GUI and strip chart screenshot.
  • 336221 10/22/10 23:38 Punjabi ROC screens
  • 336220 10/22/10 23:38 Punjabi RC and scaler screens
  • 336219 10/22/10 23:21 ZAFAR AHMED Hall C is testing feedback
  • 336216 10/22/10 22:50 Punjabi ROC screens
  • 336215 10/22/10 22:49 Punjabi RC and scaler screens
  • 336212 10/22/10 22:01 Punjabi roc sreens
  • 336211 10/22/10 22:00 Punjabi RC and scalers
  • 336210 10/22/10 21:41 ZAFAR AHMED PbF Calorimeter ADC readback
  • 336205 10/22/10 21:08 Punjabi run control, scalers screens
  • 336204 10/22/10 21:06 Punjabi Roc 17 and Roc 14 screens
  • 336203 10/22/10 20:40 gaskell Status and Runplan for remainder of swing and owl shift
  • 336201 10/22/10 20:29 Carlos jmysql is back
  • 336198 10/22/10 20:02 Nuruzzaman Target moved to LH2, LOOP 2 15 cm.
  • 336197 10/22/10 19:56 Carlos Pedestal update
  • 336196 10/22/10 19:55 ZAFAR AHMED HARP scan
  • 336195 10/22/10 19:28 Carlos jmysql server down
  • 336194 10/22/10 19:25 efuchey weird cavity behavior
  • 336193 10/22/10 18:02 hyde Deadtime study vs rate
  • 336192 10/22/10 17:52 A. Camsonne DVCS DAQ test ROC18
  • 336185 10/22/10 17:07 gaskell Beam permit
  • 336184 10/22/10 17:04 gaskell New kinematics: Kin 3, low energy
  • 336173 10/22/10 16:37 jroche pion rejector signal when HV labelled PRL1 are off
  • 336169 10/22/10 16:25 hyde clock scalers in replay
  • 336168 10/22/10 16:16 jroche, cshl Pion rejector: database detector map match reality
  • 336167 10/22/10 16:06 A. Camsonne VME ADC
  • 336165 10/22/10 16:04 Saha Shift Summary
  • 336150 10/22/10 14:20 ellie ADCs look worse after pedestal update (run 7737)
  • 336149 10/22/10 14:16 J. Zhang Moller magnet is full
  • 336148 10/22/10 13:55 Carlos Pedestals update
  • 336145 10/22/10 10:34 ellie (Run 7737) CaloAmp17 ADC looks off and has noisy TDCs
  • 336144 10/22/10 10:29 Lindgren Target charts
  • 336143 10/22/10 10:17 lindgren target photo
  • 336142 10/22/10 08:38 gaskell Status update and plan for the day
  • 336137 10/22/10 08:00 horn Owl shift summary
  • 336132 10/22/10 06:41 horn Beam back
  • 336130 10/22/10 06:31 horn Beam down for beam studies preparation
  • 336126 10/22/10 05:07 ent calo TDCs back
  • 336125 10/22/10 04:58 Carlos online gui
  • 336123 10/22/10 04:49 horn Rebooted calo HV gui
  • 336122 10/22/10 04:35 ent many calo TDC channels missing
  • 336119 10/22/10 04:18 ent problems analyzing full file
  • 336118 10/22/10 04:12 ent dead time study - runs
  • 336099 10/22/10 01:20 A. Camsonne DAQ status
  • 336098 10/22/10 01:06 A. Camsonne Run 7720 with calorimeter read out
  • 336097 10/22/10 01:05 Markowitz target status
  • 336094 10/22/10 00:57 horn Starting dead time study
  • 336092 10/22/10 00:48 horn Calorimeter HV on for test run
  • 336091 10/22/10 00:47 horn Change to nominal configuration
  • 336090 10/22/10 00:41 Markowitz more cernox L2 oscillations
  • 336086 10/22/10 00:03 Markowitz move target to LH2. 15 cm
  • 336085 10/22/10 00:02 Solvignon Swing shift summary
  • Thursday

  • 336082 10/21/10 23:50 Solvignon Spot++
  • 336080 10/21/10 23:36 cshloka target moved to Dummy, 15cm
  • 336078 10/21/10 23:08 cshloka fluctuation in cernox 2 R
  • 336075 10/21/10 22:55 Solvignon LHRS at 2.539GeV/c
  • 336073 10/21/10 22:41 A. Camsonne Pion rejector focal plane y amplitude distribution
  • 336072 10/21/10 22:36 A. Camsonne TDCs look ok now in T5
  • 336069 10/21/10 22:24 cshloka target moved to carbon
  • 336066 10/21/10 22:09 Solvignon Spot++
  • 336065 10/21/10 21:56 cshloka target moved to Dummy, 15cm
  • 336062 10/21/10 21:30 Solvignon LHRS at 2.591GeV/c
  • 336060 10/21/10 21:10 cshloka target moved to Loop 2, 15cm
  • 336058 10/21/10 20:58 Solvignon spot++: now raster pattern looks fine at 2x2
  • 336056 10/21/10 20:52 Nuruzzaman Coda crashed during begining of the run 7707.
  • 336054 10/21/10 20:45 cshloka fluctuation in cernox_2_R still exists,
  • 336051 10/21/10 20:36 Solvignon Weird raster pattern
  • 336048 10/21/10 20:26 cshloka fluctuation in cernox_2_R temperature
  • 336047 10/21/10 20:24 cshloka Target moved to Dummy, 15cm
  • 336046 10/21/10 20:18 Solvignon Cerenkov 5,6 and 7 a bit weak
  • 336044 10/21/10 19:51 cshloka target moved to Carbon
  • 336043 10/21/10 19:50 Solvignon Start LHRS delta scan: P0=2.643 GeV/c
  • 336041 10/21/10 19:44 gaskell Setting up for DIS delta scan
  • 336040 10/21/10 19:39 A. Camsonne DVCS Logic module
  • 336038 10/21/10 19:26 Carlos Calorimeter HV off
  • 336037 10/21/10 19:25 A. Camsonne Timing spectra for T3 only run
  • 336035 10/21/10 19:13 cshloka Fluctuation in cernox_2_R value
  • 336032 10/21/10 19:02 A. Camsonne ARS firmware version E10 downloaded
  • 336031 10/21/10 19:01 cshloka target status
  • 336026 10/21/10 17:55 efuchey Compton rates recheck
  • 336019 10/21/10 17:29 A. Camsonne Run 7694 Cerenkov
  • 336018 10/21/10 17:29 hyde DVCS Trig scalers
  • 336017 10/21/10 17:28 A. Camsonne Timing plots 7694 s2m TDCs
  • 336016 10/21/10 17:27 A. Camsonne Weird timing distribution in data - Run 7694
  • 336013 10/21/10 17:12 hyde DVCS Trig scalers
  • 336010 10/21/10 17:00 hyde DVCS Trig scalers
  • 336009 10/21/10 16:51 gaskell Pockels cell OFF last night
  • 336006 10/21/10 16:19 cshloka target snapshot at the start of shift
  • 336005 10/21/10 16:13 Solvignon Gas shed airflow alarm
  • 336003 10/21/10 16:01 efuchey Day shift summary 10/21/2010
  • 336000 10/21/10 15:36 A. Camsonne Run 7694 no validation run
  • 335998 10/21/10 15:32 efuchey master IOC reboot from MCC
  • 335991 10/21/10 15:21 A. Camsonne Potential good run with E09 Run 7692
  • 335990 10/21/10 15:16 A. Camsonne Readout file for ROC14 and ROC17 modified to use soft reset
  • 335989 10/21/10 15:05 A. Camsonne Oops it was run 7689
  • 335988 10/21/10 15:05 A. Camsonne Time out in read out increased Run 7688
  • 335985 10/21/10 14:51 A. Camsonne Loading ARS firmware E09 for testing
  • 335983 10/21/10 14:08 Hisham VME4: warning
  • 335981 10/21/10 13:58 Hisham Run# 7686
  • 335980 10/21/10 13:39 efuchey Compton rates for today
  • 335979 10/21/10 13:34 Hisham vme4--hatsv4-port-7 warning
  • 335977 10/21/10 13:18 Hisham Target moved to LH2, 15cm
  • 335976 10/21/10 13:12 efuchey Beam recovery procedure
  • 335975 10/21/10 12:55 Hisham Target moved to empty
  • 335974 10/21/10 12:51 hyde dvcs scalar problems
  • 335973 10/21/10 12:38 efuchey Back to beam permit
  • 335972 10/21/10 12:37 A. Camsonne DVCS DAQ work
  • 335960 10/21/10 12:11 Magali Lemo INPUT and OUTPUT of the claorimeter trigger
  • 335956 10/21/10 12:08 hyde Trigger modifications
  • 335944 10/21/10 11:31 efuchey Hall A tools screenshot
  • 335942 10/21/10 11:28 gaskell Runplan for Thursday Day --> Swing
  • 335938 10/21/10 11:00 R. Michaels attempt to get scaler server on DVCS rocs
  • 335937 10/21/10 10:43 efuchey Beam available again
  • 335936 10/21/10 10:41 efuchey Alarm from DVCS HV
  • 335935 10/21/10 10:24 R. Michaels DVCSTopBot put at top of list
  • 335934 10/21/10 10:22 J. Zhang Moller magnet has been filled
  • 335933 10/21/10 09:23 hisham Hall in controlled access
  • 335932 10/21/10 09:14 Hisham Target status snapshot
  • 335931 10/21/10 09:04 Hisham ROC 14, ROC 17
  • 335930 10/21/10 09:02 A. Camsonne Calorimeter missing channels recovered
  • 335928 10/21/10 08:25 A. Camsonne Power cycled DVCS crates
  • 335927 10/21/10 08:19 Carlos Pedestal shift in boards
  • 335923 10/21/10 08:03 horn Owl shift summary
  • 335922 10/21/10 07:51 ent end-of-run-7664 scaler screen
  • 335921 10/21/10 07:50 ent end-of-run-7664 ROC screens
  • 335920 10/21/10 07:27 A. Camsonne Waveforms block 35
  • 335918 10/21/10 07:12 A. Camsonne Missing calorimeter channels 35 and 38
  • 335915 10/21/10 07:07 A. Camsonne Run 7663 with TrigDVCS configuration
  • 335913 10/21/10 07:00 ent scaler screen run 7662
  • 335912 10/21/10 07:00 ent ROC screens
  • 335911 10/21/10 06:51 ent missing calo channels, snapshots
  • 335910 10/21/10 06:35 ent calo channels missing, runs 7660 etc.
  • 335907 10/21/10 06:16 ent end-of-run-7661 scaler screen
  • 335906 10/21/10 06:15 ent end-of-run-7661 ROC screens
  • 335902 10/21/10 04:45 horn Configuration change to DVCS production
  • 335901 10/21/10 04:29 ent boiling test runs
  • Wednesday

  • 335856 10/20/10 23:58 ZAFAR AHMED Wednesday 10-20-2010 Swing Shift Summary
  • 335851 10/20/10 23:22 Carlos ARS firmware version E09
  • 335848 10/20/10 22:27 Nuruzzaman ROC 14 and 17 errors during run 7634.
  • 335847 10/20/10 22:12 A. Camsonne Added dvcs3 flag to TrigDVCS configuration
  • 335845 10/20/10 22:00 A. Camsonne ROC18 crashed
  • 335843 10/20/10 21:51 A. Camsonne Run with Calorimeter Trigger crate run 7633
  • 335841 10/20/10 21:36 A. Camsonne DVCS Calorimeter data event dump
  • 335833 10/20/10 19:46 A. Camsonne Reverted to ARS version E08
  • 335831 10/20/10 19:31 Nuruzzaman Rebooted ROCs and Restarted CODA
  • 335830 10/20/10 19:19 gaskell Runplan for Wednesday Swing --> Thursday Day
  • 335827 10/20/10 19:04 Nuruzzaman Beam Current changed to 2 uA for run # 7625.
  • 335826 10/20/10 19:01 Nuruzzaman ROC 14 and 17 Screens during end of the run 7624.
  • 335825 10/20/10 19:01 Nuruzzaman Scaler and event rate during end of the run 7624.
  • 335823 10/20/10 18:27 Nuruzzaman ROC 14 and 17 Screens during end of the run 7623.
  • 335822 10/20/10 18:26 Nuruzzaman Scaler and event rate during end of the run 7623.
  • 335820 10/20/10 17:53 A. Camsonne Removed partial clear starting run 7623
  • 335815 10/20/10 17:04 efuchey rates variation
  • 335814 10/20/10 16:56 A. Camsonne DVCS T5 trigger with and without Cerenkov
  • 335812 10/20/10 16:42 efuchey Compton rates for today
  • 335809 10/20/10 16:39 A. Camsonne Run 7620 with version E09
  • 335808 10/20/10 16:38 A. Camsonne FPGA programmers on ARS
  • 335806 10/20/10 16:17 hyde DVCS trig outputs
  • 335805 10/20/10 16:15 A. Camsonne ARS firmware version E09 downloaded
  • 335804 10/20/10 16:01 Saha Shift Summary
  • 335793 10/20/10 14:11 hyde S1 S2 retiming
  • 335778 10/20/10 13:09 Kiadtisak Target moved to LH2
  • 335777 10/20/10 13:04 Saha Harp Scans after beam recovery
  • 335776 10/20/10 13:00 R. Michaels disk quota problems on adaq
  • 335772 10/20/10 11:36 gaskell Change of kinematics
  • 335769 10/20/10 11:08 hyde T1 input
  • 335763 10/20/10 10:50 Fansler Ice on Moeller lines
  • 335762 10/20/10 10:38 A. Camsonne Changed scaler header in datastream
  • 335761 10/20/10 10:37 hyde S0/T2 input
  • 335760 10/20/10 10:16 hyde T2 double pulsing
  • 335759 10/20/10 10:06 J. Zhang Moller magnet has been filled
  • 335758 10/20/10 09:54 jpchen Summary of cryotarget events overnight and what TO should know
  • 335756 10/20/10 07:36 gaskell Status and Controlled Access
  • 335755 10/20/10 07:33 hisham Hall in controlled access
  • 335754 10/20/10 07:31 Hisham Owl shift summary
  • 335753 10/20/10 07:11 Hisham Controlled access
  • 335752 10/20/10 07:03 Hisham Harp scan & BPM
  • 335751 10/20/10 06:26 florian itar target issue
  • 335750 10/20/10 05:49 Hisham Harp scan
  • 335749 10/20/10 05:05 Hisham Target update
  • 335747 10/20/10 04:38 Pawel run 7600 rates
  • 335746 10/20/10 04:37 Pawel run 7600 rocs
  • 335745 10/20/10 04:35 Hisham Target alarm
  • 335742 10/20/10 03:29 Pawel run 7599 rocs
  • 335741 10/20/10 03:28 Pawel run 7599 rates (end)
  • 335740 10/20/10 03:26 florian itar target problem report
  • 335739 10/20/10 03:12 florian itar alarm report
  • 335732 10/20/10 02:26 Pawel run 7599 rates
  • 335730 10/20/10 02:21 Pawel run 7598 rocs
  • 335728 10/20/10 01:24 Pawel run 7598 rates
  • 335727 10/20/10 01:16 Pawel run 7597 rocs
  • 335724 10/20/10 00:12 Solvignon Swing shift summary
  • Tuesday

  • 335722 10/19/10 23:55 Nuruzzaman ROC 14 and 17 Screens during end of the run 7596.
  • 335721 10/19/10 23:55 Nuruzzaman Scaler and event rate during end of the run 7596.
  • 335720 10/19/10 23:54 mcanan swing target screens
  • 335718 10/19/10 23:13 Nuruzzaman CODA showed error when I tried to end the run 7595.
  • 335717 10/19/10 23:07 Nuruzzaman ROC 14 and 17 Screens during end of the run 7595.
  • 335716 10/19/10 23:06 Nuruzzaman Scaler and event rate during end of the run 7595.
  • 335715 10/19/10 22:58 mcanan L2 Power is zero
  • 335712 10/19/10 22:28 mcanan L2 Power fluctuates
  • 335711 10/19/10 22:24 mcanan Cryo Temps (green and red) dropped
  • 335710 10/19/10 22:21 Nuruzzaman ROC 14 and 17 Screens during end of the run 7594.
  • 335709 10/19/10 22:21 Nuruzzaman Scaler and event rate during end of the run 7594.
  • 335707 10/19/10 21:35 Solvignon Kinematics changed to P0=2.406GeV/c, Theta=37.0
  • 335705 10/19/10 21:22 Nuruzzaman ROC 14 and 17 Screens during end of the run 7593.
  • 335704 10/19/10 21:21 Nuruzzaman Scaler and event rate during end of the run 7593.
  • 335701 10/19/10 21:01 Nuruzzaman ROC 14 and 17 Screens during end of the run 7592.
  • 335700 10/19/10 21:00 Nuruzzaman Scaler and event rate during end of the run 7592.
  • 335699 10/19/10 20:52 efuchey Table scan attempt
  • 335696 10/19/10 20:25 Nuruzzaman ROC 14 and 17 Screens.
  • 335695 10/19/10 20:24 Nuruzzaman Scaler and event rate during end of the run 7591.
  • 335692 10/19/10 19:44 Nuruzzaman Scaler Rate after Run # 7590.
  • 335691 10/19/10 19:43 Nuruzzaman ROC 14 and 17 Error.
  • 335690 10/19/10 19:34 gaskell Runplan update for Tuesday Swing --> Wednesday Owl
  • 335689 10/19/10 19:27 Nuruzzaman Run # 7590: BPM A and B.
  • 335682 10/19/10 18:45 Nuruzzaman Scaler Rate after Run # 7589.
  • 335681 10/19/10 18:44 Nuruzzaman ROC 14 and 17 error after run 7589
  • 335680 10/19/10 18:20 efuchey Rates for tonight
  • 335679 10/19/10 18:17 Nuruzzaman BPM A and B.
  • 335678 10/19/10 18:09 Nuruzzaman Analizer showing the following error.
  • 335675 10/19/10 17:39 Nuruzzaman Problem in Calorimeter ADC channel 163 and 167 ?
  • 335674 10/19/10 17:21 Nuruzzaman Error message in ROC 14 and 17
  • 335673 10/19/10 17:04 R. Michaels scalers in ROC14 and 17
  • 335671 10/19/10 16:54 Solvignon Beam back
  • 335669 10/19/10 16:19 A. Camsonne ROC14 triggers
  • 335667 10/19/10 16:04 A. Camsonne Scaler in ROC4
  • 335663 10/19/10 15:52 Hisham Day shift summary
  • 335660 10/19/10 15:11 A. Camsonne ARS in ROC14 not accepting triggers
  • 335656 10/19/10 15:04 hisham Controlled access
  • 335647 10/19/10 14:02 Hisham Empty FIFO, ROC14,Run7574
  • 335646 10/19/10 13:53 Kiadtisak Target moved to LH2
  • 335645 10/19/10 13:44 Hisham Beam Recovery
  • 335644 10/19/10 13:13 hyde DVCS DAQ deadtime study
  • 335643 10/19/10 12:58 HIsham Beam update
  • 335639 10/19/10 12:33 A. Camsonne Scalers added to VME readout
  • 335638 10/19/10 11:57 gaskell Update and plan for Day and Swing shift
  • 335637 10/19/10 11:50 hyde S0 HV
  • 335636 10/19/10 11:46 Kiadtisak Target moved to empty
  • 335635 10/19/10 11:46 Hisham Beam recovery procedure
  • 335634 10/19/10 11:46 Kiadtisak Target moved to empty
  • 335633 10/19/10 11:45 hyde S0 S2m scalers in top ARS VME crate
  • 335632 10/19/10 11:42 A. Camsonne ARS readout to reset the FIFO at every event again
  • 335631 10/19/10 11:28 A. Camsonne Clocks on scalers
  • 335625 10/19/10 10:59 J. Zhang Moller magnet has been filled
  • 335618 10/19/10 10:46 A. Camsonne Upper level Left HRS hut scope on hallavme12
  • 335617 10/19/10 10:38 A. Camsonne DVCS crates scalers assignment
  • 335607 10/19/10 08:55 Carlos ROCs at the end of run 7557
  • 335606 10/19/10 08:31 jroche Owl shift summary
  • 335604 10/19/10 08:25 jroche Hall in controlled access
  • 335603 10/19/10 08:11 jroche N2 level in Q3 really low
  • 335601 10/19/10 07:59 jroche Turn on pion rejector 1 back on
  • 335600 10/19/10 07:40 Pawel run 7557 roc17
  • 335598 10/19/10 07:26 yez a plot this shows the problem of pion rejectors
  • 335597 10/19/10 07:24 Pawel run 7557 roc14
  • 335596 10/19/10 07:17 Pawel run 7556 errors
  • 335595 10/19/10 07:13 jroche pion rejector (HV off for a test)
  • 335594 10/19/10 07:11 Pawel run 7555 roc17
  • 335592 10/19/10 06:58 Pawel run 7554 errors
  • 335589 10/19/10 06:41 jroche roc 14 and 17 at the end of run 7553
  • 335588 10/19/10 06:25 Pawel run 7553 roc 14 error
  • 335587 10/19/10 06:00 Pawel run 7553 roc17 error
  • 335586 10/19/10 05:56 Pawel coda and raw rates at 2 uA
  • 335584 10/19/10 05:48 Pawel coda and raw rates @ 3 uA
  • 335583 10/19/10 05:35 Pawel run 7551 errors
  • 335579 10/19/10 05:00 Pawel run 7551 roc17
  • 335577 10/19/10 04:50 yez target screen 2
  • 335576 10/19/10 04:49 yez target screen 1
  • 335569 10/19/10 04:21 jroche that was 4 uA
  • 335563 10/19/10 03:54 Carlos Coda and raw rates for 3uA
  • 335561 10/19/10 03:28 yez target moved to LH2 15cm in Loop2
  • 335560 10/19/10 03:22 yez target moved to BeO
  • 335559 10/19/10 03:17 jroche it's freezing in here !
  • 335558 10/19/10 03:06 Jack NMR scope signal
  • 335557 10/19/10 02:54 jroche Compton dipole problem: decided to bypass the chicane all together
  • 335551 10/19/10 02:35 jroche Status of Hall A quite misleading
  • 335550 10/19/10 02:03 efuchey QWP scan for today
  • 335549 10/19/10 01:52 jroche difficult beam recovery.
  • 335545 10/19/10 00:10 jroche NMR gui screen shot
  • 335542 10/19/10 00:00 Nuruzzaman Target GUI and stripchart screenshot at end of swing shift.
  • Monday

  • 335541 10/18/10 23:59 efuchey redid QWP scan analysis from 10-16
  • 335540 10/18/10 23:49 efuchey Results of QWP scan from 10/16/2010
  • 335539 10/18/10 23:42 BoZhao Swing shift summary
  • 335537 10/18/10 23:35 BoZhao dipole NMR signal
  • 335536 10/18/10 23:23 efuchey Results of QWP scan from 10/15/2010
  • 335534 10/18/10 23:07 Jack (Techs) LCW pressure regulator for Arc/BSY feed from Hall A is shot
  • 335533 10/18/10 23:06 A. Camsonne ARS board resetting FIFO after each event
  • 335532 10/18/10 22:56 Jack (Techs) Bogie controls issue
  • 335531 10/18/10 22:48 gaskell Runplan for Monday swing to Tuesday day
  • 335530 10/18/10 22:44 Jack Magnets status
  • 335529 10/18/10 22:44 A. Camsonne reset fan board moved to ROC14
  • 335512 10/18/10 21:42 BoZhao LHRS magnets down
  • 335500 10/18/10 19:39 gaskell LCW update
  • 335490 10/18/10 16:08 Hisham Correction: BCM calibration
  • 335488 10/18/10 16:04 Nuruzzaman Target GUI and strip chart screenshot at the begining of the swing shift.
  • 335487 10/18/10 16:02 ellie Day Shift Summary
  • 335486 10/18/10 15:59 A. Camsonne 2 ARS board surveyed and take to the CODA DAQ lab
  • 335485 10/18/10 15:28 O.Glamazdin Moller magnet has been filled
  • 335484 10/18/10 15:07 jpchen cryotarget: restoring to normal running condition
  • 335483 10/18/10 14:56 ellie LHRS Q1 flow giving yellow, occasionally red, alerts
  • 335482 10/18/10 14:31 Hisham BCM Calibration
  • 335481 10/18/10 14:04 Jack (Techs) Resets for the Bogies controls
  • 335473 10/18/10 12:41 hyde deadtime study
  • 335470 10/18/10 12:35 hyde DVCS Cosmic HV Off
  • 335469 10/18/10 12:28 Carlos Owl shift summary
  • 335463 10/18/10 11:55 jroche DVCS2 Target informations
  • 335462 10/18/10 11:39 A. Camsonne output modified
  • 335461 10/18/10 11:33 ellie HV Change for S2m L/R 12-15, Channel 8
  • 335459 10/18/10 11:31 hyde S0 HV
  • 335453 10/18/10 10:49 gaskell Status update
  • 335451 10/18/10 10:21 A. Camsonne DVCS calorimeter paddle on for cosmics data taking - Run 7500
  • 335450 10/18/10 10:02 hyde S1.and.S2 timing into T3: Correction
  • 335448 10/18/10 08:52 A. Camsonne DVCS DAQ meeting summary
  • 335446 10/18/10 07:55 jmammei target status at start of shift
  • 335444 10/18/10 06:44 yez target moved to Empty
  • 335443 10/18/10 06:04 yez target screen 2
  • 335442 10/18/10 06:03 yez target screen 1
  • 335437 10/18/10 03:31 Carlos Bogies troubles (TECHS)
  • 335436 10/18/10 03:28 Carlos Elastic setting 2
  • 335429 10/18/10 02:17 Pawel Nadel- run 7496
  • 335426 10/18/10 01:57 Pawel Nadel- run 7495
  • 335424 10/18/10 01:35 Pawel Nadel- run 7494
  • 335423 10/18/10 01:11 Carlos ARS sync problem
  • 335422 10/18/10 00:57 Pawel Nadel- run 7493
  • 335420 10/18/10 00:29 Carlos Run 7491
  • 335416 10/18/10 00:05 Carlos fifo 7490
  • 335415 10/18/10 00:03 LeRose Shift Summary (Swing)
  • Sunday

  • 335413 10/17/10 23:43 Abdurahim Loop 2 temperature sensor readout fluctuation
  • 335412 10/17/10 23:37 Abdurahim temperature sensor readout fluctuation
  • 335411 10/17/10 23:30 Ramesh fifo from 7489
  • 335407 10/17/10 23:04 Ramesh Run 7487 stopped due to fifo appearance
  • 335406 10/17/10 23:03 Abdurahim Compton
  • 335404 10/17/10 22:42 Ramesh ROC17 stopped run 7486
  • 335401 10/17/10 22:14 Ramesh Scalar rate during run 7486 @ 3 uA
  • 335400 10/17/10 22:08 LeRose Beam is back
  • 335399 10/17/10 22:07 LeRose Quad Cycling
  • 335398 10/17/10 22:06 Abdurahim Target Snapshot
  • 335397 10/17/10 22:05 Ramesh Run 7485 mostly a cosmic run
  • 335396 10/17/10 21:30 LeRose No Beam!
  • 335395 10/17/10 20:47 Carlos Calorimeter HV updated
  • 335392 10/17/10 20:25 Ramesh anti-correlation of momentum and phi angle
  • 335391 10/17/10 20:20 Abdurahim Target -> LH2 (Loop2, 15 cm)
  • 335390 10/17/10 20:18 LeRose Beam is back
  • 335389 10/17/10 20:07 Abdurahim Target -> Empty
  • 335388 10/17/10 20:04 Abdurahim Target screenshot
  • 335387 10/17/10 19:57 LeRose Spectrometer Settings
  • 335386 10/17/10 19:48 LeRose Controlled Access
  • 335379 10/17/10 18:20 Carlos Sync test run
  • 335377 10/17/10 18:08 Abdurahim Cryogen Flow Snapshots
  • 335376 10/17/10 18:07 Abdurahim Target Loops Snapshots
  • 335375 10/17/10 18:05 Abdurahim Target Screenshot
  • 335374 10/17/10 17:43 Abdurahim Target -> H2 (Loop 2, 15 cm)
  • 335373 10/17/10 17:42 A. Camsonne Clock switched
  • 335372 10/17/10 17:33 A. Camsonne ARS setup to use validation signal
  • 335371 10/17/10 17:30 Abdurahim Target -> BeO
  • 335370 10/17/10 17:19 LeRose NMR Lock
  • 335369 10/17/10 17:12 gaskell Run plan for Sunday swing and Monday owl
  • 335368 10/17/10 17:04 A. Camsonne Compton photon rates background at 4 KHz/uA
  • 335367 10/17/10 16:40 A. Camsonne Timeout in ROC14 and ROC17 set to 300
  • 335366 10/17/10 16:39 LeRose NMR Lock
  • 335365 10/17/10 16:31 Abdurahim Target -> Empty
  • 335362 10/17/10 15:59 jmammei Day Shift summary
  • 335361 10/17/10 15:52 cshloka loop 2 change in temp in cernox_2_R
  • 335360 10/17/10 15:44 cshloka empty histos in some calo TDCs
  • 335359 10/17/10 15:41 cshloka empty histos for calo adcs
  • 335358 10/17/10 15:35 cshloka no counts in some calo ADCs and TDCs
  • 335353 10/17/10 15:02 A. Camsonne Run test pulser 7474
  • 335348 10/17/10 14:37 cshloka target status
  • 335344 10/17/10 14:01 cshloka empty fifo
  • 335343 10/17/10 13:49 jmammei MOLLER refill complete
  • 335342 10/17/10 13:41 J. Zhang Moller magnet has been filled
  • 335339 10/17/10 13:00 jmammei survey complete; Hall in Controlled access
  • 335338 10/17/10 12:38 cshloka no counts in some calo TDCs
  • 335337 10/17/10 12:34 cshloka no counts in some calo ADCs
  • 335336 10/17/10 12:28 gaskell Called for Controlled Access
  • 335333 10/17/10 11:30 cshloka target status
  • 335332 10/17/10 11:26 A. Camsonne Run 7463 - test calorimeter trigger read out
  • 335331 10/17/10 11:22 A. Camsonne removing DVCS root files on disk
  • 335329 10/17/10 11:17 A. Camsonne /adaql4/work1 full
  • 335328 10/17/10 11:11 A. Camsonne adaql7 NFS issues
  • 335325 10/17/10 10:58 cshloka empty fifo for roc 14 and 17 at the end of run# 7461
  • 335319 10/17/10 10:39 gaskell Status update
  • 335318 10/17/10 10:00 cshloka empty fifo in roc 14 and 17
  • 335317 10/17/10 09:57 cshloka Empty Fifo error in ROC 17
  • 335316 10/17/10 09:55 cshloka ion rejector sum plots are blank for prl1.
  • 335315 10/17/10 09:39 jmammei changed the energy in the db_run.dat
  • 335314 10/17/10 09:32 jmammei p0=2.607; LHRS @ 34.760 degrees
  • 335311 10/17/10 09:09 A. Camsonne DVCS clock in TDC seen now
  • 335310 10/17/10 08:57 brads Beam energy a bit unstable
  • 335307 10/17/10 08:32 cshloka target status
  • 335306 10/17/10 08:17 efuchey Q1 + D alarms
  • 335305 10/17/10 08:09 efuchey 10/17/10 Owl Shift summary
  • 335303 10/17/10 07:41 A. Camsonne Requested access for HRS move back to elastic kinematics
  • 335302 10/17/10 07:40 A. Camsonne No signal seen on D.dvcsClk
  • 335298 10/17/10 07:26 A. Camsonne DVCS DAQ tests
  • 335296 10/17/10 06:23 D.Wang target status
  • 335281 10/17/10 05:01 D.Wang target moved to Loop2, 15cm
  • 335258 10/17/10 02:35 D.Wang target moved to Carbon
  • 335257 10/17/10 02:31 efuchey Hall A tools snapshot
  • 335254 10/17/10 02:19 D.Wang target moved to Loop2, 15cm
  • 335249 10/17/10 01:59 D.Wang target moved to Carbon
  • 335247 10/17/10 01:48 D.Wang target status
  • 335244 10/17/10 01:40 D.Wang target moved to Loop2, 15cm
  • 335243 10/17/10 01:30 efuchey Second harpscan
  • 335242 10/17/10 01:22 D.Wang target moved to empty
  • 335241 10/17/10 01:20 efuchey Dipole crash
  • 335240 10/17/10 01:18 efuchey Harp scan
  • 335239 10/17/10 01:01 D.Wang target moved to BeO
  • 335237 10/17/10 00:57 D.Wang spot++
  • Saturday

  • 335229 10/16/10 23:53 Ramesh Shift Summary
  • 335226 10/16/10 23:14 A. Camsonne Added clock in D class db_D.dat
  • 335225 10/16/10 22:56 A. Camsonne Clock on TDC for sync check
  • 335224 10/16/10 22:53 brads Logbook now has a preview option
  • 335223 10/16/10 22:25 Ramesh hall in controlled access
  • 335222 10/16/10 22:01 ZAFAR AHMEED Empty target
  • 335221 10/16/10 21:45 brads Tools and NMR screens for 2.0 GeV/c
  • 335220 10/16/10 21:18 Ramesh Beam down for a while for an IOC reboot
  • 335219 10/16/10 21:14 Ramesh Spectrometer set for boiling study
  • 335218 10/16/10 21:13 Abdurahim L-HRS elastic momentum vs phi (run7422)
  • 335217 10/16/10 21:04 Carlos Calorimeter HV off for boiling studies
  • 335216 10/16/10 21:03 Ramesh Preparing for boiling study
  • 335215 10/16/10 20:58 Abdurahim empity fifo5 for roc15 7 roc17
  • 335213 10/16/10 20:02 Abdurahim scalers for run 7424
  • 335211 10/16/10 19:54 Abdurahim empty histos 7422
  • 335210 10/16/10 19:48 hyde T3 timing HRS S1 S2m
  • 335207 10/16/10 19:38 Abdurahim L-HRS elastic momentum vs phi (run7422)
  • 335206 10/16/10 19:36 Abdurahim scalers window
  • 335205 10/16/10 19:35 Abdurahim xwindows for roc14 and roc17
  • 335204 10/16/10 19:24 Abdurahim L HRS elastic momentum vs phi run7421
  • 335196 10/16/10 18:32 Abdurahim screen shot for roc14 & roc17 for run7421
  • 335195 10/16/10 17:48 Abdurahim Scalers
  • 335193 10/16/10 17:33 A. Camsonne Run 7421 with test with buffering
  • 335192 10/16/10 17:25 Ramesh About to get beam
  • 335191 10/16/10 16:57 brads Tools and NMR screens for 2.406 GeV/c
  • 335190 10/16/10 16:53 A. Camsonne NMR locked at 2.406 GeV/c on probe B
  • 335189 10/16/10 16:53 Ramesh NMR locked
  • 335188 10/16/10 16:45 Ramesh Left NMR locking problem
  • 335185 10/16/10 16:31 ZAFAR AHMED Target is LH2 Loop2 15cm
  • 335184 10/16/10 16:21 ZAFAR AHMED Empty target
  • 335182 10/16/10 16:07 jmammei Day shift run summary
  • 335181 10/16/10 16:05 jmammei set LHRS to 37.000 degrees
  • 335180 10/16/10 15:59 chench HRS elastic protons
  • 335179 10/16/10 15:55 chench Run7420 has same empty Calo ADC and TDC histograms as run7419
  • 335177 10/16/10 15:47 chench screen shot for xwindow of Roc14 and roc17
  • 335176 10/16/10 15:10 chench Left HRS momentum vs phi run7419
  • 335175 10/16/10 15:02 chench Run7418,7419,7420 have same prescale set, same LHRS angle and momentum
  • 335174 10/16/10 15:00 chench Empty histogram for calo ADC and TDC in run7419
  • 335173 10/16/10 14:51 brads re: spurious gas sensor alarm?
  • 335170 10/16/10 14:42 chench Empty fifos roc14 and roc17 at the end run of 7419
  • 335169 10/16/10 14:00 chench Elastic (HRS) p vs phi
  • 335162 10/16/10 12:45 chench following figures for the scalers of Cer
  • 335160 10/16/10 12:40 chench run7418 scalers
  • 335158 10/16/10 12:37 pzhu fluctuation
  • 335156 10/16/10 12:11 chench run7417 has the same prescale and configures(leftHrs) with run7416
  • 335153 10/16/10 11:51 chench HRS elastic protons
  • 335152 10/16/10 11:39 chench following figure for scalers
  • 335151 10/16/10 11:36 chench Empty fifos for roc17
  • 335150 10/16/10 11:33 chench scalers for run7416
  • 335148 10/16/10 11:26 jmammei resuming run plan
  • 335147 10/16/10 11:22 pzhu Target moved to Loop2,15cm
  • 335146 10/16/10 11:09 jmammei MCC wants to go to 50uA; empty target
  • 335145 10/16/10 11:04 pzhu target moved to empty
  • 335144 10/16/10 11:04 jmammei MCC sending tune beam
  • 335138 10/16/10 10:46 pzhu D2Pressures increase
  • 335137 10/16/10 10:31 jmammei beam update
  • 335136 10/16/10 10:19 pzhu h2pressures increase
  • 335135 10/16/10 10:06 jmammei spurious gas sensor alarm?
  • 335134 10/16/10 09:42 jmammei NMR locked - thanks Brad!
  • 335133 10/16/10 09:28 brads LHRS dipole at 2.607 and NMR locked (HowTo instructions)
  • 335132 10/16/10 09:26 pzhu stable cryoPressures
  • 335131 10/16/10 09:11 pzhu moved target to Loop2,15cm
  • 335130 10/16/10 09:10 chench The number of events shown on coda keeping to flip all the time
  • 335129 10/16/10 08:54 pzhu fluctuation
  • 335128 10/16/10 08:51 jmammei beam down ~15 minutes
  • 335125 10/16/10 08:46 jmammei Q2, 3 momentum increased by 2%
  • 335124 10/16/10 08:45 chench spot++ for run7415
  • 335123 10/16/10 08:37 adaq moved target to BeO
  • 335121 10/16/10 08:25 chench The number of events for run7414 is flipping between two numbers
  • 335120 10/16/10 08:17 hyde Owl shift summary
  • 335119 10/16/10 08:07 jmammei NMR won't lock at 2.607 GeV
  • 335117 10/16/10 07:29 hyde HRS motion controls
  • 335116 10/16/10 06:36 hyde calo DC ADC
  • 335114 10/16/10 06:24 hyde spot++
  • 335112 10/16/10 05:34 D.Wang Target status
  • 335111 10/16/10 05:32 cshloka mpty histograms for calo TDC 129, 147 and 151
  • 335110 10/16/10 05:30 cshloka empty histograms in calo ADC 129, 147 and 151
  • 335107 10/16/10 05:14 hyde Accel down
  • 335106 10/16/10 04:52 cshloka run 7409; calo TDC 179-193 and 194-207
  • 335105 10/16/10 04:51 cshloka run 7409; calo TDC 120-143 and 144-167
  • 335104 10/16/10 04:50 cshloka run 7409; calo TDC 72-95 and 96-119
  • 335103 10/16/10 04:49 cshloka run 7409; calo TDC 24-47 and 48-71
  • 335102 10/16/10 04:47 cshloka run 7409; calo ADC 194-207 and TDC 0-23
  • 335101 10/16/10 04:46 cshloka run 7409; calo ADC 144-167 and 179-193
  • 335100 10/16/10 04:44 cshloka run 7409; calo ADC 96-119 and 120-143
  • 335099 10/16/10 04:43 cshloka run 7409; calo adc 48-71 and 72-95
  • 335098 10/16/10 04:42 cshloka calo adc 0-23 and 24-47
  • 335097 10/16/10 04:22 correction
  • 335096 10/16/10 04:16 cshloka run number for empty fifo
  • 335095 10/16/10 04:12 cshloka Empty Fifo
  • 335092 10/16/10 04:08 hyde ARS firmware
  • 335090 10/16/10 04:06 A. Camsonne Logic module FPGA loaded with 256 us dead time for valid and 1 us deadtime for stop
  • 335089 10/16/10 03:58 hyde HRS elastic protons
  • 335086 10/16/10 03:18 A. Camsonne DVCS crls files modified
  • 335085 10/16/10 03:05 hyde spot++ run 7409
  • 335084 10/16/10 02:57 hyde empty fifo
  • 335076 10/16/10 02:38 hyde Calo HV on
  • 335075 10/16/10 02:31 hyde cerenkov scalers
  • 335073 10/16/10 02:30 hyde scaler s2
  • 335072 10/16/10 02:29 A. Camsonne Calorimer HV on 400 V
  • 335071 10/16/10 02:29 hyde Scaler S1
  • 335070 10/16/10 02:28 hyde scalers on LH2
  • 335068 10/16/10 02:25 D.Wang target moved to Loo2, 15cm
  • 335066 10/16/10 02:16 cshloka moved target to BeO
  • 335065 10/16/10 02:08 hyde bPM
  • 335060 10/16/10 01:39 D.Wang target status
  • 335057 10/16/10 01:02 A. Camsonne DVCS DAQ work
  • 335054 10/16/10 00:56 hyde bbscope/ HRS
  • 335053 10/16/10 00:54 hyde S1 S2 TDC
  • 335049 10/16/10 00:46 hyde 1024 Hz Clk Double Pulsing
  • 335048 10/16/10 00:45 hyde Cerenkov TDC Threshold
  • Friday

  • 335037 10/15/10 23:56 BoZhao Swing Shift Summary
  • 335036 10/15/10 23:23 J. Zhang Moller magnet has been filled
  • 335035 10/15/10 23:18 BoZhao Raster is turned off
  • 335034 10/15/10 23:06 O.Glamazdin Moller measurements result
  • 335033 10/15/10 21:50 hyde BCM
  • 335032 10/15/10 21:31 hyde BCM calibration
  • 335031 10/15/10 20:38 O.Glamazdin Moller magnet status
  • 335030 10/15/10 20:32 O.Glamazdin Moller SC magnet quench
  • 335029 10/15/10 20:10 hyde BCM
  • 335023 10/15/10 18:46 Abdurahim Compton Cavity QWP Scan Procedures
  • 335022 10/15/10 18:07 Kalyan Follow-up on lumi test
  • 335019 10/15/10 16:48 brads How to Reset the HRS bogie IOC
  • 335018 10/15/10 16:36 efuchey QWP scan for 10/15/2010
  • 335016 10/15/10 15:54 X. Zheng Day shift summary
  • 335015 10/15/10 15:48 brads Run Plan for Friday 15 DAY -> Saturday 16 DAY
  • 335014 10/15/10 15:19 Saha Hall A Arc Energy Measurement
  • 335013 10/15/10 15:09 X. Zheng Arc done, now Moller
  • 335012 10/15/10 14:10 efuchey change of Compton threshold
  • 335011 10/15/10 14:10 efuchey change of Compton threshold
  • 335010 10/15/10 14:09 A. Camsonne scope in LHRS plugged on network bbscope1
  • 335009 10/15/10 13:21 X. Zheng Arc energy measurement
  • 335008 10/15/10 12:42 brads DVCS Calorimeter HV off
  • 335006 10/15/10 11:50 A. Camsonne DVCS calorimeter HV ON for cosmics
  • 335005 10/15/10 11:46 A. Camsonne Logic module FPGA loaded with 256 us dead time for stop
  • 335004 10/15/10 11:14 ellie Run 7388, CaloAmp6 ADC looks wonky. Lots of TDCs have low signal:noise
  • 335003 10/15/10 11:08 ellie Pion Rejector analysis plots blank
  • 335002 10/15/10 11:01 A. Camsonne BCM log file
  • 334996 10/15/10 10:44 O.Glamazdin Moller magnet has been filled
  • 334995 10/15/10 10:33 kelleher Target motion -- empty
  • 334992 10/15/10 09:32 X. Zheng Current plan
  • 334990 10/15/10 09:13 X. Zheng Left HRS moved to 33.290 degrees
  • 334989 10/15/10 09:04 R. Michaels re: helicity in scalers
  • 334988 10/15/10 08:53 jroche new Helicity board firmware installed
  • 334985 10/15/10 08:04 jroche Owl shift summary
  • 334982 10/15/10 06:55 florian itar calo TDC 194-207
  • 334981 10/15/10 06:54 florian itar calo TDC 144-167 and 179-193
  • 334980 10/15/10 06:52 florian itar calo TDC 96-119 and 120-143
  • 334979 10/15/10 06:51 florian itar calo TDC 48-71 and 72-95
  • 334978 10/15/10 06:50 florian itar calo TDC 0-23 and 24-47
  • 334977 10/15/10 06:49 florian itar calo ADC 194-207
  • 334976 10/15/10 06:48 florian itar calo ADC 144-167 and 179-193
  • 334975 10/15/10 06:47 florian itar calo ADC 96-119 and 120-143
  • 334974 10/15/10 06:45 florian itar calo ADC 48-71 and 72-95
  • 334973 10/15/10 06:44 florian itar calo ADC 0-23 et 24-47
  • 334971 10/15/10 05:49 florian correction
  • 334968 10/15/10 05:06 florian itar current information
  • 334967 10/15/10 05:03 florian counting rates current information
  • 334966 10/15/10 05:00 yez target screen - cryo
  • 334965 10/15/10 04:59 yez target screen -- temp & press.
  • 334964 10/15/10 04:55 florian itar counting rates elastic calibration calorimeter
  • 334963 10/15/10 04:54 florian itar counting rates elastic calibration calorimeter
  • 334962 10/15/10 04:36 jroche Left bogie being really difficult (TECHS)
  • 334960 10/15/10 03:41 jroche LHRS set for 2.607 GeV
  • 334959 10/15/10 03:20 jroche shirt access to check on ladders
  • 334952 10/15/10 02:48 yez Sieve Partten Analysis -- See holes for LH2 but not for Carbon foils
  • 334949 10/15/10 02:13 florian itar counting rates elastic calibration calorimeter
  • 334948 10/15/10 02:12 florian itar counting rates elastic calibration calorimeter
  • 334939 10/15/10 00:00 Solvignon Swing shift summary
  • Thursday

  • 334936 10/14/10 23:53 A. Camsonne Calorimeter HV on - Run 7377 1 uA
  • 334934 10/14/10 23:36 efuchey Compton rates became dramatic
  • 334926 10/14/10 22:42 A. Camsonne Calorimeter HV on - Run 7371
  • 334922 10/14/10 22:32 R. Michaels expected rates = measured rate, I think.
  • 334921 10/14/10 21:51 R. Michaels lack of holes for T3
  • 334920 10/14/10 21:38 R. Michaels LH2 rates, run 7369
  • 334919 10/14/10 21:36 R. Michaels study to analyze trigger efficiency
  • 334918 10/14/10 21:22 Solvignon Spot++
  • 334916 10/14/10 21:13 mcanan Target Motion
  • 334911 10/14/10 20:36 hyde s0 tdc in raw data
  • 334910 10/14/10 20:34 mcanan Target Motion
  • 334909 10/14/10 20:30 efuchey Cavity scan
  • 334908 10/14/10 19:54 mcanan Target Motion
  • 334906 10/14/10 19:45 Kalyan Summary of lumi test runs
  • 334899 10/14/10 19:24 hyde S1 S2 trigger timing in TDC
  • 334898 10/14/10 19:23 Kalyan Run numbers of Lumi test A
  • 334893 10/14/10 19:04 efuchey Update for Compton rates
  • 334891 10/14/10 19:03 Kalyan HV setting for lumi test B
  • 334889 10/14/10 18:55 efuchey lost readout of compton:rate_G
  • 334874 10/14/10 18:15 efuchey Compton rates
  • 334872 10/14/10 18:11 Kalyan HV setting for Lumi test A
  • 334871 10/14/10 18:05 brads Injector matching/tuning was completed before the BCM calibration started
  • 334870 10/14/10 18:01 brads Notes on the BCM calibration procedure
  • 334869 10/14/10 17:56 mcanan Target Motion
  • 334868 10/14/10 17:43 brads re: Left HRS NMR -- multiple probes available
  • 334867 10/14/10 17:36 hyde EDTM in S2m TDC
  • 334865 10/14/10 17:09 gomez Left HRS NMR -- multiple probes available
  • 334862 10/14/10 16:46 Solvignon BCM calibration: 1st part
  • 334856 10/14/10 15:49 R. Michaels network work finished, MSS copying resumed
  • 334855 10/14/10 15:36 Carlos Left sieve is OUT
  • 334854 10/14/10 15:25 Jack Magnets use LCW - They will need to be ramped down and turned off
  • 334853 10/14/10 14:38 brads LHRS NMR tricks and tips
  • 334851 10/14/10 14:18 Magali Which ARS board in which crate
  • 334850 10/14/10 14:05 J. Zhang Moller magnet has been filled.
  • 334849 10/14/10 13:57 folts/tech sieve position
  • 334848 10/14/10 13:31 brads Update on LHRS sieve and NMR status
  • 334839 10/14/10 12:37 brads Hall will be down on Monday, 0700 for LCW work
  • 334832 10/14/10 11:56 hyde HRS-DVCS trigger timing
  • 334830 10/14/10 11:51 brads Run Plan for Thurs 14 DAY -> Friday 15 DAY
  • 334813 10/14/10 10:16 efuchey 1mm thick Lead filter put back
  • 334812 10/14/10 10:15 efuchey 1mm thick Lead filter put back
  • 334810 10/14/10 09:45 R. Michaels network switchover in progress
  • 334807 10/14/10 09:02 jroche owl shift summary
  • 334804 10/14/10 08:05 jroche Scinitllator timing: things look better
  • 334801 10/14/10 07:33 yez LH2 Loop temperature was a little big unstable but fine now
  • 334798 10/14/10 06:34 jroche Cerenkov calibration constant: new values
  • 334797 10/14/10 06:33 jroche Cerenkov calibration constant: new values
  • 334791 10/14/10 05:38 adaq scaler data
  • 334789 10/14/10 05:26 jroche Compton: hall C to 10 uA makes rates A-ok for our Comptom
  • 334787 10/14/10 04:51 jroche LHRS to second elastic setting for calorimeter calibration.
  • 334786 10/14/10 04:50 jroche Trouble moving LHRS from angle 34.76 to 37.00 (TECHS)
  • 334785 10/14/10 04:48 yez target screen -- cryo
  • 334784 10/14/10 04:47 yez target screen -- temp and press
  • 334783 10/14/10 04:28 jroche Trouble getting the LHRS Dipole NMR to lock (TECHS)
  • 334774 10/14/10 02:17 jroche Compton rates got higher: aborted Compton data taking
  • 334769 10/14/10 01:09 jroche left LHRS angle display in General tools.
  • Wednesday

  • 334761 10/13/10 23:40 jroche LHRS angle in Hall A general Tool is way off
  • 334757 10/13/10 23:03 ZAFAR AHMED Shift summary Wednesday October 13
  • 334747 10/13/10 22:33 A. Camsonne Low rate ARS pedestal run 7303
  • 334746 10/13/10 22:25 A. Camsonne coda EB seg faulting at end of run
  • 334745 10/13/10 22:23 A. Camsonne coda EB seg faulting at end of run
  • 334739 10/13/10 21:59 A. Camsonne DVCS ARS Calorimeter pedestal run 7296
  • 334736 10/13/10 21:32 mcanan loop 2 temperature (green)
  • 334733 10/13/10 20:49 A. Camsonne Access for DVCS
  • 334732 10/13/10 20:35 mcanan Loop 2 temperature change hal2_Cernox_2_R
  • 334731 10/13/10 20:31 M Friend Compton rates
  • 334728 10/13/10 20:06 Hisham HRS Singles
  • 334724 10/13/10 20:00 ZAFAR AHMED Beam permit
  • 334723 10/13/10 19:39 M Friend Compton photon beamline
  • 334722 10/13/10 19:14 ZAFAR AHMED Controlled access
  • 334721 10/13/10 18:57 ZAFAR AHMED Sive is out
  • 334719 10/13/10 18:49 ZAFAR AHMED Beam current is 5uA
  • 334711 10/13/10 18:38 brads Run Plan for Wed 13 SWING -> Thurs 14 DAY
  • 334709 10/13/10 18:19 ZAFAR AHMED Hall A tools
  • 334708 10/13/10 18:16 ZAFAR AHMED Target is LH2
  • 334707 10/13/10 18:16 mcanan Target Motion
  • 334706 10/13/10 18:07 mcanan Target Motion
  • 334705 10/13/10 18:04 ZAFAR AHMED Spot++
  • 334703 10/13/10 18:03 ZAFAR AHMED Beam is back.
  • 334691 10/13/10 17:08 ZAFAR AHMED Beam Update
  • 334689 10/13/10 16:45 ZAFAR AHMED Beam update
  • 334687 10/13/10 16:34 brads Target GUI Labels are now correct
  • 334685 10/13/10 16:19 mcanan Target Screens
  • 334684 10/13/10 16:19 ZAFAR AHMED Beam permit
  • 334682 10/13/10 16:03 saha Shift Summary
  • 334680 10/13/10 15:46 hyde Changes to trigger retiming
  • 334679 10/13/10 15:44 A. Camsonne pulser in ARS 0x8
  • 334678 10/13/10 15:30 A. Camsonne Left HRS set to 2.606 GeV/c for elastic
  • 334671 10/13/10 14:12 J. Zhang Moller has been refilled
  • 334670 10/13/10 14:12 J. Zhang Moller has been refilled
  • 334668 10/13/10 14:06 A. Camsonne iocha14 rebooted by itself ?
  • 334663 10/13/10 13:54 hsmith Left Dipole polarity change
  • 334662 10/13/10 13:45 R. Michaels problems with trigger downloading (and remedy)
  • 334657 10/13/10 13:36 Carlos HRS & calorimeter at elastic positions
  • 334649 10/13/10 12:59 kelleher loop 2 temperature (sensor?)
  • 334644 10/13/10 12:43 efuchey Calorimeter RG 213 cables tested
  • 334643 10/13/10 12:01 atrig Trigger-Download
  • 334642 10/13/10 11:58 atrig Trigger-Download
  • 334641 10/13/10 11:55 atrig Trigger-Download
  • 334640 10/13/10 11:51 atrig Trigger-Download
  • 334639 10/13/10 11:50 atrig Trigger-Download
  • 334638 10/13/10 11:49 atrig Trigger-Download
  • 334637 10/13/10 11:47 atrig Trigger-Download
  • 334636 10/13/10 11:46 atrig Trigger-Download
  • 334635 10/13/10 11:42 atrig Trigger-Download
  • 334634 10/13/10 11:36 kelleher Hall A status
  • 334633 10/13/10 11:20 atrig Trigger-Download
  • 334632 10/13/10 11:08 atrig Trigger-Download
  • 334622 10/13/10 10:45 kelleher target motion -- empty
  • 334620 10/13/10 10:34 hyde S1 HV
  • 334619 10/13/10 10:31 hyde S2m timing
  • 334609 10/13/10 10:01 kelleher target motion -- LH2 15cm
  • 334608 10/13/10 09:57 Saha Arc Scanner Test
  • 334605 10/13/10 09:29 kelleher target motion -- optics target
  • 334603 10/13/10 08:54 A. Camsonne Wrong order in FPGA chain
  • 334600 10/13/10 08:04 jroche owl shift summary (better formating )
  • 334599 10/13/10 08:02 jroche owl shift summary
  • 334595 10/13/10 06:15 jroche target is LH2 not C foils
  • 334594 10/13/10 06:11 jroche Scintillator timing (S2 only in the trigger)
  • 334592 10/13/10 04:59 yez target move to LH2 15cm Loop2
  • 334591 10/13/10 04:44 yez target move to empty
  • 334590 10/13/10 04:12 yez target move to Loop2 LH2 15cm
  • 334589 10/13/10 04:03 yez target moves to LH2 15cm
  • 334587 10/13/10 03:28 yez target screem2 - cryo gui
  • 334586 10/13/10 03:27 yez target screen 1 - L1,L2
  • 334579 10/13/10 02:44 jroche sieve back in position
  • 334578 10/13/10 02:40 yez Cannot see Multi-foils partten with Seive Slit in, but it is clear when Sieve Slit is out
  • 334567 10/13/10 01:14 jroche sieve out
  • 334563 10/13/10 00:05 Solvignon Shift summary
  • Tuesday

  • 334559 10/12/10 22:00 mcanan Target motion
  • 334558 10/12/10 21:53 mcanan Target Motion
  • 334557 10/12/10 21:52 Solvignon Starting beam recovery procedure
  • 334556 10/12/10 21:47 mcanan Target motion
  • 334555 10/12/10 21:10 mcanan Target Screens 10_12 Swing
  • 334554 10/12/10 20:18 Solvignon Starting compton test with different beam currents
  • 334553 10/12/10 20:00 efuchey Compton Rates after tuning
  • 334552 10/12/10 19:04 Solvignon Compton tune in progress
  • 334546 10/12/10 17:45 brads Helicity flip at 943.3Hz
  • 334545 10/12/10 17:42 brads Run Plan for Tues 12 SWING -> Wed 13 DAY
  • 334543 10/12/10 17:36 Solvignon cw beam
  • 334542 10/12/10 17:31 mcanan Hall C target
  • 334541 10/12/10 17:29 mcanan Loop 2 hp power and JT Valve
  • 334540 10/12/10 17:27 brads re: Carbon foils operational current limit: 25 uA
  • 334539 10/12/10 17:23 Solvignon Beam tuning
  • 334536 10/12/10 17:07 Solvignon/Ca EPICS gateway problem
  • 334534 10/12/10 16:50 A. Camsonne DVCS DAQ access
  • 334533 10/12/10 16:49 J. Zhang Moeller magnet has been filled.
  • 334531 10/12/10 16:38 Solvignon Helicity flipping rate is now 960Hz
  • 334530 10/12/10 16:36 M Friend Compton DAQ trigger back to normal
  • 334529 10/12/10 16:34 mcanan loop 2 minor alarm
  • 334527 10/12/10 16:21 Solvignon Hall A in controlled access
  • 334520 10/12/10 11:18 Carlos BCM calibration aborted
  • 334499 10/12/10 08:44 Carlos BCM calibration setup
  • 334498 10/12/10 08:40 Carlos BCM calibration starting
  • 334497 10/12/10 08:39 X. Zheng Target moved to empty
  • 334494 10/12/10 08:15 jroche Owl shift summary
  • 334493 10/12/10 07:54 jroche changing conditions: still looking for a reason
  • 334489 10/12/10 06:34 jroche beam operation autorization
  • 334488 10/12/10 05:57 yez Graphs were all the same. Please check the most updated ones
  • 334484 10/12/10 03:33 jroche sieve in position
  • 334483 10/12/10 03:21 florian itar target
  • 334481 10/12/10 03:16 jroche Carbon foils operational current limit: 25 uA
  • 334479 10/12/10 02:57 jroche Left HRS to 3.327 GeV
  • 334473 10/12/10 02:28 jroche Q1 mistmatch fixed
  • 334472 10/12/10 02:28 jroche Q1 mistmatch fixed
  • 334471 10/12/10 02:13 jroche Q1 magnet mismatch??
  • 334469 10/12/10 01:57 A. Camsonne Trigger VME interface stuck
  • 334464 10/12/10 01:07 M Friend Compton background rates
  • 334463 10/12/10 00:59 A. Camsonne VME bus analyzer unplugged
  • 334462 10/12/10 00:40 A. Camsonne FPGA order on USB2
  • 334460 10/12/10 00:01 Jin Huang Elastic stripe, dp = -4%
  • Monday

  • 334459 10/11/10 23:57 Jin Huang Shift Summary : Oct 11 Swing
  • 334458 10/11/10 23:48 yez Proton Mass Spectrum for Elastic Data with different Center Momentum Setting
  • 334456 10/11/10 23:45 Jin Huang Hall A tool & HV @ dp = -4% elastic scan
  • 334454 10/11/10 23:43 Jin Huang Spot++ @ LH2, 2uA CW
  • 334453 10/11/10 23:40 BoZhao screen of target charts
  • 334452 10/11/10 23:40 yez Beam Spot on BeO with Raster off
  • 334451 10/11/10 23:39 BoZhao target moved to LH2@15cm
  • 334448 10/11/10 23:36 Jin Huang Spot++, 5uA CW on BeO
  • 334446 10/11/10 23:28 Jin Spot++ with tune beam on LH2
  • 334445 10/11/10 23:24 BoZhao target moved to BeO
  • 334443 10/11/10 23:15 BoZhao Target moved to loop2 LH2@15cm
  • 334435 10/11/10 22:11 Jin Huang Update: Beam is back
  • 334434 10/11/10 21:40 BoZhao screen of target charts
  • 334433 10/11/10 20:11 Jin Huang Update : No Beam. Asked for beam permit
  • 334432 10/11/10 19:44 A. Camsonne Pulses in ARS
  • 334430 10/11/10 19:20 Alex ARS boards in crates
  • 334429 10/11/10 19:20 Alex
  • 334428 10/11/10 19:02 A. Camsonne FPGA programmers assignement
  • 334422 10/11/10 18:29 hyde DVCS cables under door
  • 334421 10/11/10 17:55 M Friend Temporary Compton DAQ trigger change
  • 334420 10/11/10 17:52 BoZhao screen of target charts
  • 334417 10/11/10 16:41 riordan Latest update is beam at 7:00PM
  • 334410 10/11/10 15:59 BoZhao Screen of target charts
  • 334402 10/11/10 15:36 riordan Left arm crate HV off
  • 334401 10/11/10 14:59 Holmstrom Target CommunicationPrb?
  • 334400 10/11/10 14:39 Kalyan follow-up on lumi analysis
  • 334399 10/11/10 14:05 A. Camsonne New version of ARS firmware
  • 334398 10/11/10 13:31 J. Butler Moeller
  • 334397 10/11/10 13:24 Jack (Techs) Left dipole NMR stability
  • 334396 10/11/10 12:24 Kalyan Re: Correction: Lumi linearity
  • 334395 10/11/10 11:40 Kalyan Correction: Lumi linearity
  • 334389 10/11/10 10:35 A. Camsonne DVCS crates boot script modified
  • 334388 10/11/10 09:04 O.Glamazdin Moller magnet is back to operation
  • 334387 10/11/10 08:51 riordan Rebooted iochla14
  • 334386 10/11/10 08:40 riordan Gateway restored but p0 setting not reading
  • 334385 10/11/10 08:29 riordan Injector work during day shift
  • 334383 10/11/10 05:10 Hisham Owl shift summary
  • 334382 10/11/10 04:43 Hisham Owl shift summary
  • 334381 10/11/10 04:28 A. Camsonne EPICS gateway issues affect magnets regulation - giving up on beam for now
  • 334380 10/11/10 04:19 riordan Stopping for tonight - hall to controlled access
  • 334378 10/11/10 04:07 Hisham Spot++
  • 334376 10/11/10 04:04 Hisham Beam is back
  • 334374 10/11/10 03:57 Hisham NMR is locked
  • 334373 10/11/10 03:20 A. Camsonne Swing shift DAQ work
  • 334372 10/11/10 03:17 A. Camsonne Hall A Epics gateway down
  • 334371 10/11/10 02:50 riordan IOCs down
  • 334365 10/11/10 02:43 Hisham NMR is not locked
  • 334364 10/11/10 02:09 hisham Controlled acces
  • 334363 10/11/10 02:08 hisham Late entry: Beam is back
  • 334361 10/11/10 02:06 hisham Late entry: Out of the Hall
  • 334351 10/11/10 00:36 Hisham Screen snapshot: NMR
  • 334341 10/11/10 00:00 LeRose Shift Summary (Swing)
  • Sunday

  • 334334 10/10/10 23:47 jroche Vertex resolution with new t0 for vdc
  • 334323 10/10/10 23:17 jroche Cerenkov PMT1 now showing a nice pe peak
  • 334280 10/10/10 20:18 A. Camsonne DVCS bottom crate firmware update from file E06.pof
  • 334279 10/10/10 19:58 Jin Huang Target: 2nd temperature sensor on loop2 just fixed itself
  • 334250 10/10/10 18:14 L Controlled Access
  • 334249 10/10/10 17:51 riordan Run Plan for Swing 10/10 through Day 10/11
  • 334248 10/10/10 17:37 LeRose Controlled Access
  • 334247 10/10/10 17:11 Carlos ARS tests (run 7116)
  • 334246 10/10/10 17:09 Kalyan Current scan for lumi detector
  • 334245 10/10/10 17:08 Jin Huang Target -> Empty
  • 334243 10/10/10 16:53 Carlos ARS tests
  • 334242 10/10/10 16:48 A. Camsonne crl files copied in Oct-10-2010
  • 334241 10/10/10 16:46 A. Camsonne crl files copied in Sept9-2010
  • 334240 10/10/10 16:26 Jin Huang Target: Heater upper limit from 300W->500W for both loops
  • 334238 10/10/10 16:00 jmammei Day shift summary
  • 334234 10/10/10 15:43 A. Camsonne Spot++ run 7123 on LH2 5uA
  • 334233 10/10/10 15:32 Bellini Target move to LH2 15cm
  • 334232 10/10/10 15:27 A. Camsonne Raster check run 7122
  • 334229 10/10/10 14:42 jmammei raster was reset to 4x4mm^2
  • 334228 10/10/10 14:24 jmammei turned spectrometer HV on
  • 334227 10/10/10 14:23 O.Glamazdin Moller magnet
  • 334226 10/10/10 14:13 jmammei Sasha out of hall
  • 334225 10/10/10 14:08 jmammei controlled access for MOLLER
  • 334224 10/10/10 13:27 jpchen temperature spikes due to noise
  • 334220 10/10/10 12:52 jmammei Carlos came out of hall
  • 334219 10/10/10 12:52 Carlos Pulse signal plugged into lower ARS crate
  • 334218 10/10/10 12:32 jmammei Jixie came out of Hall
  • 334217 10/10/10 12:31 J. Zhang Moller magnet has been refilled
  • 334216 10/10/10 12:28 jmammei raster decreased for MOLLER
  • 334215 10/10/10 12:27 jmammei Jixie and Carlos go in for escorted access
  • 334214 10/10/10 12:13 Bellini Target screens
  • 334213 10/10/10 12:12 jmammei cryo temp drop
  • 334209 10/10/10 11:50 Bellini Transient red alarm on Loop1
  • 334208 10/10/10 11:47 Bellini Screen snapshot
  • 334206 10/10/10 11:36 A. Camsonne ROC18 crashed
  • 334205 10/10/10 11:27 jmammei calorimeter HV back off
  • 334202 10/10/10 11:14 jmammei turned off spectrometer HV
  • 334200 10/10/10 10:56 Magali run on ARS Board at the electronic Lab
  • 334199 10/10/10 10:48 jmammei preparing for MOLLER tests
  • 334193 10/10/10 10:32 Bellini Target moved to empty
  • 334192 10/10/10 09:37 Carlos Run pulser 7103
  • 334191 10/10/10 09:35 Carlos Run pulser 7103
  • 334190 10/10/10 08:58 jmammei NMR lock lost briefly - fixed itself
  • 334189 10/10/10 08:43 cshloka Cryo levels
  • 334187 10/10/10 08:00 M Friend Owl shift summary
  • 334186 10/10/10 07:07 M Friend NMR lost lock temporarily
  • 334185 10/10/10 06:26 M Friend NMR temporarily lost lock again
  • 334184 10/10/10 05:42 M Friend NMR temporarily lost lock
  • 334183 10/10/10 05:13 D.Wang target moved Loop2, 15cm
  • 334182 10/10/10 05:05 itard spot++
  • 334181 10/10/10 05:05 D.Wang target moved to BeO
  • 334179 10/10/10 04:31 D.Wang target moved to empty
  • 334175 10/10/10 03:11 D.Wang target status
  • 334167 10/10/10 02:15 M Friend Beam down
  • 334165 10/10/10 01:46 A. Camsonne DVCS DAQ status
  • 334164 10/10/10 01:34 D.Wang target status
  • 334162 10/10/10 01:31 M Friend Calorimeter HVs OFF
  • 334160 10/10/10 01:12 M Friend Hall back to beam permit
  • 334159 10/10/10 01:11 M Friend Calorimeter HVs on
  • 334158 10/10/10 01:07 A. Camsonne Calorimeter VME interface board unplugged
  • 334157 10/10/10 01:03 M Friend Hall in Controlled access
  • 334154 10/10/10 00:51 A. Camsonne Channels with pulser
  • 334152 10/10/10 00:40 M Friend Calorimeter HV GUI problem
  • 334148 10/10/10 00:24 M Friend GC ch. 1 HV to 1800V
  • 334146 10/10/10 00:01 Jin Huang Shift Summary : Oct 09 Swing
  • Saturday

  • 334145 10/09/10 23:54 jroche new T0 for the Left VDC updated in the database
  • 334144 10/09/10 23:31 A. Camsonne DVCS calorimeter HV off
  • 334143 10/09/10 23:26 A. Camsonne Typo in onlana.odef
  • 334135 10/09/10 23:00 Jin and Alex LHRS p0 -> 3.392GeV/c (-4%)
  • 334134 10/09/10 22:50 A. Camsonne NMR current setting
  • 334130 10/09/10 22:42 A. Camsonne Left HRS HV off for run 7095
  • 334129 10/09/10 22:39 A. Camsonne Calorimeter random test top crate Run 7094
  • 334128 10/09/10 22:38 Jin & Alex CaF2 cal. ADC readback
  • 334127 10/09/10 22:31 A. Camsonne Calorimeter HV changed to 400 V and turned on for random data taking on calorimeter
  • 334126 10/09/10 22:28 Jin & Alex PbF2 Cal. ADC Read back
  • 334124 10/09/10 21:33 salvatore Cherenkov
  • 334122 10/09/10 21:14 Jin Huang HV change on Gas Cer
  • 334120 10/09/10 21:09 jroche Cerenkov (partial) calibration
  • 334119 10/09/10 21:06 Jin Huang Update on run database (online replay)
  • 334118 10/09/10 20:54 Min Huang warning of haL1_cernox_3_R
  • 334117 10/09/10 20:33 Jin Huang Elastic Kinematics Check Plots
  • 334116 10/09/10 20:22 Jin Huang Beam current -> 10uA
  • 334114 10/09/10 20:06 Jin Huang Cryo flow back. LH2 stablized. Continue taking data
  • 334112 10/09/10 20:03 Jin Huang Stop beam due to drop of 15K cryo supply
  • 334111 10/09/10 20:00 Jin Huang JT valve change, beam current adjustment
  • 334106 10/09/10 18:48 A. Camsonne DVCS logic module updated
  • 334099 10/09/10 18:40 Jin Huang LHRS p0->3.327GeV/c, (+0% to elastic)
  • 334097 10/09/10 17:42 Jin Huang Open up JT valve of Loop2 (LH2) to 47.8%(+2%)
  • 334094 10/09/10 17:09 mhuang target screen shot
  • 334093 10/09/10 17:09 salvatore Cherenkov counter6
  • 334092 10/09/10 16:48 Kalyan Online lumi monitoring
  • 334090 10/09/10 16:30 A. Camsonne Gas cerenkov channel 6 no pedestal or single photoelectron peak
  • 334085 10/09/10 16:05 Carlos Day shift summary
  • 334078 10/09/10 15:32 adaq The beam current for run7082 is:1.0uA, not 0.5uA
  • 334075 10/09/10 15:18 efuchey Hall A tools scrennshots
  • 334073 10/09/10 15:09 Bellini Twin red alarm for Loop1
  • 334071 10/09/10 14:55 chench Are the missing pion rejection ADCs out of acceptance?
  • 334067 10/09/10 14:10 efuchey HRS NMR lock unstable
  • 334066 10/09/10 13:36 Kalyan Lumi HV settings for 10uA on H2 15cm (loop2)
  • 334064 10/09/10 13:18 Bellini Move back to LH2 target
  • 334061 10/09/10 12:49 efuchey Issues with beam
  • 334060 10/09/10 12:46 Bellini Moving to empty target
  • 334059 10/09/10 12:26 efuchey HRS-L momentum change
  • 334058 10/09/10 11:54 efuchey HRS-L change
  • 334057 10/09/10 11:40 O.Glamazdin Moller SC magnet has been filled
  • 334056 10/09/10 11:32 Bellini HI and HIHI limits Change in Gas Panels
  • 334055 10/09/10 10:57 riordan Going to Controlled Access
  • 334054 10/09/10 10:54 Bellini / Ca Yellow persistent alarm on LH2
  • 334048 10/09/10 10:35 Bellini Red transient alarm on L1
  • 334047 10/09/10 10:35 Magali state of the ARS test in the Electronic Lab
  • 334046 10/09/10 10:34 Magali state of the ARS test in the Electronic Lab
  • 334045 10/09/10 10:33 riordan HRS to p0 = 3.775 GeV/c
  • 334042 10/09/10 09:38 A. Camsonne Scaler rates LH2 10 uA after trigger download
  • 334041 10/09/10 09:30 A. Camsonne Trigger for Left HRS redownloaded
  • 334039 10/09/10 09:28 atrig Trigger-Download
  • 334035 10/09/10 08:57 Carlos Rates at p0=3.932 GeV sieve OUT
  • 334034 10/09/10 08:53 Bellini Target Move to LH2 15cm
  • 334033 10/09/10 08:50 Carlos Beam's back
  • 334032 10/09/10 08:40 Carlos New HRS setting
  • 334031 10/09/10 08:23 Carlos Sieve pattern
  • 334030 10/09/10 08:17 Bellini Target moved to empty
  • 334028 10/09/10 08:10 Carlos Owl shift summary
  • 334027 10/09/10 08:09 Carlos Beam off for 30 min
  • 334026 10/09/10 07:53 Carlos Email test
  • 334025 10/09/10 07:27 riordan Too many factors to consider resolution right now
  • 334022 10/09/10 06:29 Carlos Rates with sieve IN
  • 334021 10/09/10 06:25 Carlos Rates at 10uA on 15cm LH2
  • 334020 10/09/10 06:22 Carlos Elastic scattering on LH2
  • 334019 10/09/10 06:21 D.Wang target status
  • 334015 10/09/10 05:39 A. Camsonne Vacuum stil pumping in HRS ?
  • 334010 10/09/10 04:51 Carlos Raster 7x7 on LH2
  • 334007 10/09/10 04:35 Carlos spot++ on LH2
  • 334006 10/09/10 04:29 Carlos Vertex resolution
  • 334005 10/09/10 04:23 D.Wang Target status
  • 334003 10/09/10 04:15 Carlos Optics reconstruction
  • 334001 10/09/10 04:07 D.Wang Target moved to Loop2(LH2), 15cm
  • 333998 10/09/10 02:57 D.Wang Target moved to Optics
  • 333997 10/09/10 02:57 Carlos Raster 5.5x4
  • 333989 10/09/10 02:16 Carlos spot++ 2x2
  • 333988 10/09/10 02:07 Carlos Calorimeter cosmics runs
  • 333986 10/09/10 01:54 D.Wang Target moved to Carbon Hole
  • 333984 10/09/10 01:48 D.Wang target status
  • 333983 10/09/10 01:47 D.Wang target moved from empty to BeO
  • 333982 10/09/10 01:44 Carlos Calorimeter HV off
  • 333981 10/09/10 01:39 Carlos Beam ON
  • 333979 10/09/10 00:39 Carlos Sweep in progress
  • 333977 10/09/10 00:29 jpchen GUIs may still have wrong lables
  • 333976 10/09/10 00:26 jpchen target screens
  • 333974 10/09/10 00:21 jpchen cryotarget cooldown complete
  • 333972 10/09/10 00:01 jroche for c S2 ADC basically already matched
  • Friday

  • 333971 10/08/10 23:37 jpchen Instruction for Switch between LH2 (loop2) and LD2 (loop1)
  • 333970 10/08/10 23:23 A. Camsonne DVCS calorimeter cosmics run 7061
  • 333969 10/08/10 23:15 A. Camsonne DVCS calorimeter cosmics run 7061
  • 333966 10/08/10 23:06 A. Camsonne Calorimeter HV on for cosmics tests
  • 333962 10/08/10 22:52 A. Camsonne ARS Firmware E06 loaded
  • 333961 10/08/10 22:46 A. Camsonne FPGA programmers
  • 333960 10/08/10 21:53 riordan Sieve is OUT
  • 333959 10/08/10 21:47 riordan HRS to p0 = 4.050 GeV
  • 333958 10/08/10 21:07 Bo D2 pressure vs Loop1 temperature
  • 333957 10/08/10 20:40 jpchen Loop 1 is D2 and is already cold now.
  • 333956 10/08/10 19:34 Bo An alarm of D2 low Pressure
  • 333955 10/08/10 19:17 Bo screen of target temperature
  • 333949 10/08/10 18:03 A. Camsonne Laptop in HRS
  • 333947 10/08/10 17:55 riordan Plans for Tonight
  • 333946 10/08/10 17:52 jpchen cryotarget status: waiting for Qweak target cool-down with uncertainty
  • 333944 10/08/10 17:04 jroche Left TSETTLE plugged in trigger 9
  • 333943 10/08/10 17:02 jroche helicity frequency test
  • 333937 10/08/10 16:29 gomez Left HRS -- Q2/Q3 degaussed & momentum set to 3.5 GeV
  • 333935 10/08/10 16:20 andrew/tech checklist
  • 333934 10/08/10 16:08 yez target screen -- loop1,loop2
  • 333933 10/08/10 16:07 yez target screen - main screen
  • 333932 10/08/10 16:06 yez target screen - temp,press. and flow
  • 333930 10/08/10 15:32 jroche plugged a copy of TSETTLE in TS9
  • 333921 10/08/10 15:00 Carlos Calorimeter at 5m
  • 333920 10/08/10 11:34 O.Glamazdin Moller SC magnet has been filled
  • 333914 10/08/10 09:58 jpchen Qweak and Hall A target status and our target cooldown plan
  • 333911 10/08/10 07:44 Bellini Target report
  • 333905 10/08/10 00:31 Bellini screen snapshot
  • Thursday

  • 333904 10/07/10 23:49 Jin Huang Target Screen 6
  • 333903 10/07/10 23:19 Jin Huang Target Screen 5
  • 333902 10/07/10 21:35 Jin Huang Target Screen 4
  • 333901 10/07/10 21:06 riordan Run Plan for Swing 10/7 through Day 10/8, plans into next week
  • 333900 10/07/10 19:34 Jin Huang Target Screen 3
  • 333899 10/07/10 18:50 Jin Huang Target Screen 2.1
  • 333893 10/07/10 18:38 jpchen cryotargets status
  • 333892 10/07/10 18:34 Jin Huang Target Screen 2/2
  • 333891 10/07/10 18:34 Jin Huang Target Screen 1/2
  • 333890 10/07/10 17:16 J. Zhang Moller magnet is full
  • 333888 10/07/10 16:13 M Friend Compton data from Oct 1st
  • 333887 10/07/10 16:12 Andrew/tech Checklist
  • 333886 10/07/10 16:03 hsmith Q1 magnet power supply
  • 333885 10/07/10 15:53 hsmith Q2 magnet power supply
  • 333883 10/07/10 14:07 jpchen Reduced coolant flow/fan speeds
  • 333882 10/07/10 12:31 jpchen Hall C cryotarget cooldown started
  • 333881 10/07/10 10:35 A. Camsonne ARS test runs in DAQ room list
  • 333880 10/07/10 10:24 riordan Late Entry: Hall in Restricted Access
  • 333878 10/07/10 09:36 yez cryotarget: yellow arlarm shows up-> haL1_hp_power_R
  • 333877 10/07/10 09:29 yez Cryotarget RED alarm shows up again: halTC502_1_Sensor_3_R
  • 333876 10/07/10 09:13 yez cryotarget screen 2
  • 333875 10/07/10 09:13 yez cryotarget screens
  • 333874 10/07/10 08:01 Bellini Errata Corrige
  • 333873 10/07/10 06:29 Carlos Owl shift summary
  • 333872 10/07/10 06:22 Carlos Dipole ramped down
  • 333871 10/07/10 06:17 Carlos Controlled access requested
  • 333870 10/07/10 06:12 Carlos Calorimeter HV reduced
  • 333869 10/07/10 06:10 Carlos Lumi tests
  • 333867 10/07/10 06:10 Carlos DC current with beam OFF
  • 333866 10/07/10 06:08 Carlos Lumi tests
  • 333865 10/07/10 06:05 Carlos PMT at 600V DC current
  • 333861 10/07/10 05:56 brads re: Raster 4x4 on C hole (shouldn't worry about non-centered hole)
  • 333860 10/07/10 05:44 Bellini Red Alarm On Target Loop 1
  • 333859 10/07/10 05:44 Carlos 7x6 raster on C hole
  • 333857 10/07/10 05:26 Carlos Raster 4x4 on C hole
  • 333848 10/07/10 04:18 Carlos Target to Carbon hole
  • 333847 10/07/10 04:18 Carlos Moved target to BeO
  • 333846 10/07/10 04:05 riordan Ramping up Dipole to 200A
  • 333845 10/07/10 04:04 M Friend Compton tune not successful
  • 333844 10/07/10 04:00 Carlos Compton tuning stopped
  • 333843 10/07/10 03:41 Carlos Run 7041, event 23000
  • 333842 10/07/10 03:27 Carlos Calorimeter DAQ analysis
  • 333841 10/07/10 00:42 Carlos Beam back
  • 333839 10/07/10 00:08 adaq Target Screen Shot
  • Wednesday

  • 333838 10/06/10 23:21 M Friend Compton tuning begun
  • 333837 10/06/10 21:39 M Friend Beam problems
  • 333836 10/06/10 21:10 A. Camsonne Issues to be checked from run 7041
  • 333835 10/06/10 20:15 M Friend Hall A energy stripchart
  • 333834 10/06/10 19:34 efuchey PbF2 Calorimeter HV Off
  • 333832 10/06/10 19:03 Solvignon Beam tuning starting
  • 333831 10/06/10 19:00 Jack (Techs) Magnet testing status
  • 333830 10/06/10 18:54 A. Camsonne Event Dump Run 7041 event 88
  • 333829 10/06/10 18:45 A. Camsonne ARS Cosmics run
  • 333822 10/06/10 17:52 A. Camsonne crl directory
  • 333820 10/06/10 17:45 A. Camsonne HRS HV off for tests
  • 333818 10/06/10 16:41 riordan Run Plan for Swing 10/6 through Day 10/7
  • 333816 10/06/10 16:01 andrew/tech checklist
  • 333815 10/06/10 14:36 R. Michaels end-run scripts
  • 333812 10/06/10 14:15 yez cryo warning sound is gone and the presure is going back.
  • 333811 10/06/10 14:15 techs Q1
  • 333809 10/06/10 14:08 yez cryotarget low presure warning!
  • 333806 10/06/10 13:39 O.Glamazdin Moller SC magnet has been filled
  • 333805 10/06/10 13:08 hsmith Q2 Over temperature faults
  • 333804 10/06/10 11:56 jpchen LD2 pressure
  • 333793 10/06/10 11:14 yez L2Temps went high
  • 333791 10/06/10 10:59 hyde Compton eDet noise
  • 333790 10/06/10 10:55 yez cryotarget screens-Loop1&Loop2
  • 333789 10/06/10 10:54 yez cryotarget control screen
  • 333782 10/06/10 10:27 yez Cryotarget screen
  • 333781 10/06/10 09:46 yez hacsbc2 automatically rebooted
  • 333765 10/06/10 00:13 jroche, cshl BPMs have been calibrated
  • 333764 10/06/10 00:10 adaq cryotarget screens
  • 333763 10/06/10 00:06 jpchen cryotarget cooldown completed
  • Tuesday

  • 333754 10/05/10 21:54 Magali Magne Timing on bus reduction for ARS
  • 333752 10/05/10 20:24 jpchen switched to cold-return
  • 333746 10/05/10 18:19 M Friend Compton FADC DAQ clock change
  • 333745 10/05/10 18:06 jroche Pedestals for s1,s2,cer and pir updated
  • 333744 10/05/10 17:37 jpchen J-T valve for loop 3
  • 333743 10/05/10 16:45 R. Michaels epicsLogger script split in two
  • 333741 10/05/10 16:23 riordan Update to run plan
  • 333740 10/05/10 16:18 jpchen Re-start target cool-down
  • 333735 10/05/10 15:40 Carlos Calorimeter anode current check
  • 333734 10/05/10 15:31 J. Zhang Moller solenoid has been refilled
  • 333732 10/05/10 14:42 jroche Calorimeter DC monitor GUI now reliable
  • 333731 10/05/10 12:51 A. Camsonne Top crate first registers read comparison
  • 333730 10/05/10 12:51 A. Camsonne Top crate first registers read comparison
  • 333725 10/05/10 12:15 A. Camsonne DVCS calorimeter HV back on
  • 333724 10/05/10 12:15 hisham ADC synch checker cables are installed
  • 333721 10/05/10 12:00 jpchen cryotarget cooldown status
  • 333714 10/05/10 11:39 A. Camsonne ARS Firmware D26 loaded
  • 333713 10/05/10 11:39 riordan Run Plan for Swing 10/5 through Day 10/6
  • 333708 10/05/10 11:23 M Friend Moved Compton FADC DAQ BCM
  • 333706 10/05/10 11:20 hyde HRS Left Cerenkov Threshold
  • 333692 10/05/10 09:55 Hisham Helicity cables
  • 333688 10/05/10 06:04 riordan Requested Restricted Access
  • 333687 10/05/10 04:01 Hisham Owl shift summary
  • 333685 10/05/10 03:59 Carlos BPM calibration (cont)
  • 333669 10/05/10 01:42 Carlos Calorimeter anode current
  • 333665 10/05/10 00:18 Carlos BPM calibration (cont)
  • Monday

  • 333663 10/04/10 23:53 Jin Huang Shift Summary : Oct 04 Swing
  • 333661 10/04/10 23:52 Jin Huang Target -> empty
  • 333658 10/04/10 23:29 Jin Huang Moving -> Optics target
  • 333657 10/04/10 23:18 Jin Huang Moving to BeO
  • 333655 10/04/10 23:12 Jin Huang Screen shot: Tune beam in hall
  • 333651 10/04/10 22:14 Jin Huang Tuning beam into hall
  • 333650 10/04/10 22:06 R. Michaels EPICS truncation due to ET insertions
  • 333649 10/04/10 21:44 R. Michaels re: EPICS problem : how to know if it's a problem with Podd or not
  • 333648 10/04/10 21:10 Jin Huang Update: Hall remain Controled Access due to injector issues
  • 333647 10/04/10 21:02 Jin Huang Sweep finished
  • 333646 10/04/10 20:12 Carlos Calorimeter HV off
  • 333645 10/04/10 20:01 Carlos Hall A locking up in 15min
  • 333644 10/04/10 19:56 yez L-HRS Pion Rejectors Plots for cosmic ray run6988
  • 333643 10/04/10 19:53 yez L-HRS Gas Cerenkov Plots for cosmic ray run6988
  • 333642 10/04/10 19:52 yez L-HRS Scin Plots for cosmic ray run6988 (Same plots, but with names)
  • 333641 10/04/10 19:51 meekins Target Pump Purge
  • 333640 10/04/10 19:48 yez L-HRS Scin Plots for cosmic ray run6988
  • 333639 10/04/10 19:36 yez L-HRS VDC plots for cosmic ray run6988
  • 333633 10/04/10 18:32 jroche Pedestals need adjusting: will do that tomorrow
  • 333632 10/04/10 18:01 Carlos EPICS truncation (cont)
  • 333631 10/04/10 17:33 riordan Run Plan for Swing 10/4 through Day 10/5
  • 333628 10/04/10 16:18 Kalyan lumi channel 7 and 8
  • 333626 10/04/10 15:59 A. Camsonne ARS cosmics run 6987 firmware D24
  • 333625 10/04/10 15:46 A. Camsonne Signal tap - second file working
  • 333623 10/04/10 15:42 andrewl Cryo target pre beam checklist
  • 333621 10/04/10 15:36 A. Camsonne DVCS calorimeter cosmics scintillators paddle HV on
  • 333620 10/04/10 15:29 R. Michaels new helicity patterns (thanks, Qweak)
  • 333616 10/04/10 15:18 A. Camsonne HRS HV off for DAQ test
  • 333611 10/04/10 14:52 hisham Helicity signals
  • 333610 10/04/10 14:50 A. Camsonne Calorimeter HV on for cosmics tests
  • 333609 10/04/10 14:02 R. Michaels ADC synch checker installed
  • 333597 10/04/10 10:58 riordan Hall in Restricted Access
  • 333591 10/04/10 09:47 jmammei Day Shift Summary
  • 333587 10/04/10 06:57 Carlos Owl shift summary
  • 333586 10/04/10 06:46 Carlos 2uA on optics target
  • 333585 10/04/10 06:43 Carlos MCC will run multi-harp scan
  • 333583 10/04/10 06:07 Carlos Target back to empty
  • 333581 10/04/10 05:58 Carlos Move target to optics
  • 333579 10/04/10 05:53 Carlos Raster 7x5
  • 333577 10/04/10 05:51 Carlos Raster calibration
  • 333574 10/04/10 05:45 Carlos 3x2 raster
  • 333568 10/04/10 04:43 Carlos Raster calibration
  • 333567 10/04/10 04:43 Carlos BPM calibration runlist
  • 333566 10/04/10 04:42 Carlos BPM calibration aborted
  • 333542 10/04/10 00:02 efuchey compton rates for tonight
  • Sunday

  • 333540 10/03/10 23:58 Carlos Compton rates still extremely high
  • 333537 10/03/10 23:14 Carlos Moller commissioning finished
  • 333536 10/03/10 23:07 O.Glamazdin Moller results
  • 333535 10/03/10 23:03 O.Glamazdin Moller results
  • 333534 10/03/10 21:33 O.Glamazdin Moller solenoid quench
  • 333533 10/03/10 19:24 jroche bpms are working: need harps scan for calibration
  • 333526 10/03/10 17:18 riordan Run Plan for Swing 10/3 through Day 10/4
  • 333516 10/03/10 15:23 A. Camsonne Rough BCM calibration
  • 333505 10/03/10 14:05 A. Camsonne Spot BPM pedestals updated
  • 333504 10/03/10 14:00 A. Camsonne Spot Left DB backed up
  • 333503 10/03/10 13:53 A. Camsonne BPM pedestals from get_bpm_pedestals.C
  • 333500 10/03/10 13:46 A. Camsonne BPM pedestal runs 6950,6951
  • 333498 10/03/10 13:36 A. Camsonne Requested BPM gains set to 0 for BPM pedestal run
  • 333495 10/03/10 13:11 Magali aquisition rack status
  • 333494 10/03/10 12:51 A. Camsonne BPM offset and attenuation adjusted
  • 333493 10/03/10 12:49 A. Camsonne BPM offset and attenuation adjusted
  • 333492 10/03/10 12:30 benesch EES checked out their piece of the 4A/4B BPMs
  • 333491 10/03/10 10:57 Fansler Cryo
  • 333484 10/03/10 06:10 Carlos Going to controlled access
  • 333483 10/03/10 05:47 Carlos Owl shift summary
  • 333479 10/03/10 05:39 Carlos Done for tonight
  • 333477 10/03/10 05:38 Carlos Calorimeter anode current study
  • 333476 10/03/10 05:37 D.Wang target moved to Empty
  • 333474 10/03/10 05:32 Carlos Calorimeter with HV off
  • 333473 10/03/10 05:30 Carlos Calorimeter 3uA on optics target
  • 333470 10/03/10 05:23 D.Wang Target moved to Optics
  • 333469 10/03/10 05:22 Carlos 1 uA on optics target
  • 333468 10/03/10 05:21 Carlos Optics target in
  • 333465 10/03/10 05:14 Carlos Calorimeter at 2uA in C
  • 333464 10/03/10 05:09 Carlos Calorimeter at 1uA on C
  • 333463 10/03/10 05:01 Carlos Carbon target in
  • 333462 10/03/10 04:53 Carlos BPM studies
  • 333461 10/03/10 04:49 Carlos BPMs at 10uA
  • 333459 10/03/10 04:37 Carlos BPM signals at 2uA
  • 333446 10/03/10 03:15 Carlos BPM working at extremely low current
  • 333441 10/03/10 02:52 Carlos Compton studies aborted
  • 333437 10/03/10 02:45 efuchey Compton Rates
  • 333429 10/03/10 02:32 Carlos Beam trips
  • 333422 10/03/10 01:33 Carlos Beam restoration...
  • 333417 10/03/10 00:53 efuchey Compton Rates too high tonight
  • Saturday

  • 333405 10/02/10 23:04 O.Glamazdin Moller commissioning
  • 333401 10/02/10 22:24 Roman FADC Data Trigger Time Distribution
  • 333398 10/02/10 20:40 riordan Shift Update
  • 333397 10/02/10 18:40 Carlos Work done during controlled access
  • 333391 10/02/10 16:27 riordan Run Plan for Swing 10/2 - Day 10/3
  • 333387 10/02/10 16:06 A. Camsonne Run 6914 Top crate 592 D22
  • 333385 10/02/10 15:57 A. Camsonne ARS board 592 firmware D22 loaded
  • 333381 10/02/10 15:00 riordan beam permit for Moller
  • 333380 10/02/10 14:25 fansler LQ2
  • 333379 10/02/10 13:42 Fansler Cryo
  • 333375 10/02/10 12:10 fansler Moeller
  • 333365 10/02/10 10:22 Fansler Moeller N2
  • 333364 10/02/10 10:18 fansler Cryo
  • 333358 10/02/10 00:37 riordan Requested Controlled Access
  • 333357 10/02/10 00:09 Xiaohui Zhan Xiaohui Zhan
  • Friday

  • 333356 10/01/10 23:24 efuchey Compton Scanning
  • 333355 10/01/10 22:37 Carlos BMP wires 3 and 4 showing wrong readings
  • 333354 10/01/10 21:23 Kalyan Lumi test runs 6908 6907
  • 333352 10/01/10 20:51 Kalyan lumi test runs
  • 333351 10/01/10 20:48 Kalyan Target empty
  • 333348 10/01/10 20:33 Kalyan Lumi high voltage screen shot
  • 333347 10/01/10 20:31 Kalyan Target move to optics
  • 333346 10/01/10 20:30 Kalyan target moved to carbon
  • 333343 10/01/10 19:52 A. Camsonne One missing histogram for pion rejector
  • 333342 10/01/10 19:51 riordan BeO Spot on Target for BPM (0,0)
  • 333341 10/01/10 19:47 Carlos BPM readout problem
  • 333337 10/01/10 19:45 riordan Target to Single Carbon Foil
  • 333332 10/01/10 19:39 riordan Target to BeO
  • 333322 10/01/10 18:23 jroche,hisha Width of discriminated cerenkov signals
  • 333321 10/01/10 17:44 riordan RC Experiment Summary, Oct 1, 2010
  • 333317 10/01/10 17:02 riordan Target Motion Operational
  • 333316 10/01/10 16:51 jpchen solid target can be used for checkout with low current
  • 333308 10/01/10 13:30 Techs Cryo Target Pre Beam Checklist
  • 333301 10/01/10 11:39 hyde bpm in ROC 4
  • 333300 10/01/10 11:32 hyde roc 4 slot 23 gate cable
  • 333298 10/01/10 11:03 Carlos 2nd DVCS HV crate moved to port 6 of dvcstlab1
  • 333291 10/01/10 10:17 M Friend Compton EPICS variable change
  • 333289 10/01/10 10:05 Carlos Lumi "pedestal" run 6892
  • 333288 10/01/10 09:22 R. Michaels roc4 missing a gate AGAIN
  • 333286 10/01/10 07:00 riordan Hall to Restricted Access
  • 333280 10/01/10 01:55 efuchey Compton rates: beam in chicane 2
  • 333279 10/01/10 01:44 adaq Controlled access requested
  • 333278 10/01/10 01:32 efuchey compton rates: beam in chicane
  • 333277 10/01/10 00:48 Carlos Beam established
  • 333276 10/01/10 00:28 Carlos Beam through chicane